### v1

# MedFact: Benchmarking the Fact-Checking Capabilities of Large Language Models on Chinese Medical Texts

Abstract
The increasing deployment of Large Language Models (LLMs) in healthcare necessitates a rigorous evaluation of their factual reliability. However, existing benchmarks are often limited by narrow domains of data, failing to capture the complexity of real-world medical information. To address this critical gap, we introduce MedFact, a new and challenging benchmark for Chinese medical fact-checking. MedFact comprises 2,116 expert-annotated instances curated from diverse real-world texts, spanning 13 medical specialties, 8 fine-grained error types, 4 writing styles, and multiple difficulty levels. Its construction employs a hybrid AI-human framework where iterative expert feedback refines an AI-driven, multi-criteria filtering process, ensuring both high data quality and difficulty. We conduct a comprehensive evaluation of 20 leading LLMs, benchmarking their performance on veracity classification and error localization against a human expert baseline. Our results reveal that while models can often determine if a text contains an error, precisely localizing it remains a substantial challenge, with even top-performing models falling short of human performance. Furthermore, our analysis uncovers a frequent “over-criticism” phenomenon, a tendency for models to misidentify correct information as erroneous, which is exacerbated by advanced reasoning techniques such as multi-agent collaboration and inference-time scaling. By highlighting these critical challenges for deploying LLMs in medical applications, MedFact provides a robust resource to drive the development of more factually reliable and medically aware models. Content warning: This paper contains content that may be inappropriate or offensive.
footnotetext: Corresponding authors.
Our data and code: https://github.com/ivy3h/MedFact.
Introduction
Large Language Models (LLMs) (OpenAI 2024a; DeepSeek-AI 2025a; He et al. 2025; Tang et al. 2025), renowned for their versatile capabilities in natural language processing (NLP), are increasingly being applied to real-world medical applications (Huang et al. 2024; Wu et al. 2025), such as clinical decision support (Rao et al. 2023), patient assessment (Zack et al. 2024), diagnosis (Kuroiwa et al. 2023), and text classification (Zhang et al. 2022).
While various benchmarks exist to assess the medical capabilities of LLMs, they typically focus on tasks like medical question answering (QA)(Jin et al. 2021) and relation extraction(Guan et al. 2020). However, the crucial issue of their factual reliability in the medical domain remains underexplored. This gap is particularly alarming given the potential for severe harm from medical misinformation (Wang and Togher 2024).
Existing efforts to bridge this gap are limited. For instance, VeriFact (Chung et al. 2025) evaluates fact-checking on synthetic, LLM-generated clinical text. MEDEC (Abacha et al. 2025) is restricted to error detection in clinical notes. By focusing on synthetic data or a single text genre, these benchmarks fail to capture the diversity of medical information encountered in the real world. This narrow focus precludes a comprehensive assessment of LLM fact-checking capabilities and highlights the need for a benchmark covering a broader range of realistic medical content.
To address these limitations, we introduce MedFact, a challenging benchmark for Chinese medical fact-checking, which is built upon three core principles: rigorous design, broad coverage, and uncontaminated evaluation.
  • Rigorously Designed Pipeline: We construct MedFact through a human-in-the-loop pipeline that integrates large-scale AI filtering with fine-grained annotation by medical professionals. We also employ hard-case mining to systematically retain challenging instances designed to probe the limits of current LLMs.
  • Broad and Realistic Coverage: MedFact is curated from diverse real-world texts such as medical encyclopedias. It comprises 2,116 expert-annotated medical texts that span 13 medical specialties, 8 fine-grained error types, 4 writing styles, and multiple difficulty levels.
  • Uncontaminated Data: MedFact is constructed from proprietary texts under copyright-compliant agreements, making it highly unlikely to be part of the common pre-training corpora for LLMs. This helps ensure a fairer assessment of their true fact-checking capabilities.
We benchmark 20 leading proprietary and open-source LLMs on MedFact across two distinct tasks: veracity classification (VC) and the more demanding error localization (EL). Our experiments reveal that even state-of-the-art models exhibit significant limitations, with performance falling substantially short of human expert levels. Moreover, while retrieval-augmented generation (RAG) can enhance performance, we find that other strategies, such as multi-agent collaboration and inference-time budget forcing, induce an overly critical bias, causing models to misclassify correct texts as incorrect frequently. This finding underscores the critical gap between current LLM capabilities and the requirements for safe deployment in high-stakes medical applications. Our main contributions are:
  • We introduce MedFact, a novel benchmark for Chinese medical fact-checking, distinguished by its rigorous construction, broad coverage of real-world medical content, and its design to mitigate training data contamination.
  • We present a comprehensive evaluation of 20 LLMs, revealing a significant performance gap relative to human experts on both veracity classification and the precise localization of factual errors.
  • We conduct in-depth analyses of model behavior, examining the impact of advanced strategies like multi-agent collaboration and inference-time scaling. Our findings reveal an “over-criticism” phenomenon, and demonstrate that MedFact remains a challenge even for state-of-the-art models using sophisticated inference techniques.
Figure 1: An overview of the data construction pipeline and the components of MedFact.
Related Works
LLMs in Medical Domains
The significant capabilities of Large Language Models (LLMs) across many disciplines have prompted the exploration of their applications in the medical domain (Huang et al. 2024; Chen et al. 2025). Medprompt (Nori et al. 2023) enhances LLMs’ performance on medical benchmarks by combining self-generated Chain-of-Thought (CoT) (Wei et al. 2022), few-shot learning with dynamic example selection (Liu et al. 2022), and self-consistency (Wang et al. 2023b). Other developments include HuaTuo (Wang et al. 2023a), a LLaMA-based model fine-tuned on Chinese medical instruction data; HuatuoGPT (Zhang et al. 2023), a model specialized for medical consultation, fine-tuned on responses distilled from ChatGPT and real-world data; and PediatricsGPT (Yang et al. 2024), a model developed via a systematic training pipeline for pediatric and general medical applications. These advances highlight a growing interest in adapting LLMs to the challenges of medical reasoning and the specialized requirements of real-world practice.
Benchmarking Medical Knowledge in LLMs
Numerous benchmarks have been developed to evaluate the medical knowledge of LLMs (Nguyen et al. 2023; Kweon et al. 2024; Abacha et al. 2025; Wu et al. 2025). MedQA (Jin et al. 2021) is a multiple-choice QA dataset derived from the United States Medical Licensing Examinations to assess the clinical knowledge of LLMs. More recently, HealthBench (Arora et al. 2025) is a benchmark of realistic health conversations evaluated with custom, physician-created rubrics. MedXpertQA (Zuo et al. 2025) evaluates expert-level medical knowledge using both textual and multimodal tasks.
While some benchmarks for medical fact-checking exist, they often focus on narrow contexts. VeriFact (Chung et al. 2025) verifies synthetic text against structured Electronic Health Records using a retrieval-augmented generation (RAG) and LLM-as-a-Judge paradigm. MEDEC (Abacha et al. 2025) provides a benchmark for error detection and correction, but is confined to clinical notes. Relying on synthetic data or a single text genre, these benchmarks do not capture the diverse contexts of medical information encountered in the real world, limiting their utility for assessing how well models generalize. Consequently, a research gap remains for a benchmark that can robustly evaluate fact-checking across a wide array of realistic medical texts.
Benchmarking Factuality in LLMs
Recent research has produced several benchmarks to evaluate factuality in LLMs. For instance, HaluEval (Li et al. 2023) focuses on hallucination detection through a large-scale collection of automatically generated and human-annotated examples. SimpleQA (Wei et al. 2024) measures short-form factuality using single-answer questions adversarially collected to challenge state-of-the-art models like GPT-4. While these benchmarks provide valuable tools for assessing general-domain factuality, they lack the domain-specific nuances required for high-stakes fields like medicine, where misinformation can have severe public health consequences (Wang and Togher 2024). Addressing this challenge requires a solution that bridges two critical gaps: the lack of domain-specific context in general factuality benchmarks and the lack of real-world diversity in existing medical fact-checking benchmarks. To this end, we introduce MedFact, a new benchmark explicitly designed to evaluate the fact-checking capabilities of LLMs on a comprehensive and diverse collection of realistic medical texts.
MedFact
Data Collection
The construction pipeline for MedFact is illustrated in Figure 1. We began with a corpus of 27,116 medical texts secured through a copyright-compliant data-sharing agreement with a commercial partner. The source materials are primarily drawn from the partner’s internal medical encyclopedia, supplemented with content from its medical consultant platforms, including question-answering pages and user forum discussions. As these sources have not been previously released to the public, this approach substantially mitigates the risk of training data contamination.
Data Curation
Multi-criteria Filtering We defined a suite of seven leading LLMs: GPT-4o (OpenAI 2024b), GPT-4.5 (OpenAI 2025a), o1 (OpenAI 2024c), Gemini 2.5 Pro (DeepMind 2025), DeepSeek-R1 (DeepSeek-AI 2025a), Claude 3.7 Sonnet (Anthropic 2025), and Doubao-Seed-1.6-thinking (ByteDance 2025). This model suite formed an ensemble that, via majority vote, filtered the initial corpus by discarding texts identified as overly simplistic (e.g., lacking medical knowledge), esoteric (e.g., highly specialized content), or malformed (e.g., missing a clear subject or context).
This automated process was refined through a three-round, iterative human-in-the-loop feedback mechanism inspired by Wu et al. (2024). In each round, medical experts reviewed a 5% sample of filtered data. We then used their annotations on misclassified items to construct a more effective retrieval-augmented, few-shot prompt for the subsequent round. This prompt leveraged criterion-specific retrieval, augmenting each input with the most semantically similar expert-annotated example for each of our three filtering criteria: simplistic, esoteric, and malformed. Over three rounds, the data acceptance rate decreased from 67.69% to 37.00%, and finally to 23.62%. The resulting filter achieved 96.40% agreement with expert judgments, reducing the candidate pool from 27,116 to 6,405 texts.
Human Annotation The filtered corpus was then annotated by a team of three fully licensed medical professionals. The annotation for each text comprised: (a) a binary veracity judgment (correct/incorrect); (b) for incorrect texts, the precise error span; and (c) a suggested correction.
Quality Enhancement and Refinement We then applied a series of automated techniques to enhance the dataset’s quality and difficulty, reducing the number of instances to 2,862. These steps included:
  • Hard-case Mining: To increase the benchmark’s ability to differentiate among competent models, we evaluated all annotated instances with our model suite and removed those that all seven models classified correctly, thereby filtering for more challenging cases.
  • Similarity Filtering: To promote topical and semantic diversity, we computed sentence embeddings for all texts using all-mpnet-base-v2 (Reimers and Gurevych 2019). For any pair of texts with a cosine similarity score exceeding a threshold of 0.9, we retained only one instance.
  • Data Augmentation: To enhance linguistic diversity and ensure privacy, we prompted a randomly selected model from our suite to paraphrase and de-identify each text under two strict constraints: (a) preserving the original medical content; and (b) removing all Personally Identifiable Information (PII).
Final Human Review In the final stage, the dataset underwent a comprehensive review by medical professionals. They verified that: (a) automated paraphrasing did not introduce errors or alter the original medical intent; (b) all PII was successfully removed; and (c) annotations remained accurate for the paraphrased texts. Any text failing these checks was either corrected or discarded. This process resulted in the final MedFact of 2,116 instances, which was balanced to contain an equal number of correct and incorrect texts.
Figure 2: An example of fact-checking performed by Claude 3.7 Sonnet. The medical error is in red, and key steps in the model’s reasoning are in blue. The original Chinese text and reasoning are presented first, followed by their English translations.
Figure 3: Distribution of the medical specialties in MedFact.
Data Statistics & Analysis
Overview MedFact contains 2,116 expert-annotated medical texts curated from real-world data. The dataset comprises 1,058 medically correct texts and 1,058 texts with a single factual error. It was constructed using a hybrid AI-human framework that combines large-scale multi-criteria filtering, hard-case mining, similarity filtering, and LLM-based augmentation with human expert review to ensure high data quality, diversity, and difficulty. Spanning 13 medical specialties, instances in MedFact are categorized by 8 fine-grained error types and multiple difficulty levels to facilitate a comprehensive evaluation of LLM fact-checking capabilities. Figure 2 presents a sample instance; additional examples are provided in the Appendix.
Data Diversity MedFact covers 13 medical specialties, with a distribution led by Internal Medicine (1,156 instances), followed by Obstetrics and Gynecology (240), Surgery (231), and ten other fields. These are grouped into four high-level domains: Internal Medicine and Chronic Care; Surgical and Procedural Care; Specialized Clinical Fields; and Public and Mental Health (Figure 3). This distribution reflects a realistic range of medical topics, consistent with the distribution of 27,116 original medical texts, enabling robust evaluation of model performance across diverse clinical contexts. Beyond its topical breadth, MedFact exhibits significant stylistic diversity. The corpus is drawn from a variety of writing styles, including formal encyclopedia articles rich in professional terminology, popular science journalism, user-generated web content from both peer-to-peer and patient-physician Q&A, and fabricated or sensationalized misinformation. This diversity in topic and writing style ensures that MedFact mirrors the complexity of real-world medical information, facilitating a robust evaluation of an LLM’s generalizable fact-checking capabilities.
Error Taxonomy Developed in consultation with medical professionals, our error taxonomy for MedFact distinguishes between two primary categories: medical and non-medical errors. Medical errors are subdivided into Conceptual, Terminological, Temporal, and Other Medical Errors, while non-medical errors are classified as Citation and Sourcing, Biased Content, General Factual, and Other Non-Medical Errors. Detailed definitions for each subcategory are available in the Appendix. As shown in Figure 1, the distribution across the 1,058 incorrect instances is heavily skewed toward medical errors (89.41%), with Conceptual errors being the most frequent subtype (52.65% of all incorrect instances). The prevalence of these medically grounded errors underscores that successful performance on this benchmark requires a deep understanding of medical principles.
Difficulty Grading Instances in MedFact are stratified into five difficulty levels. This stratification is based on the aggregated success rate of our model suite on the error localization task (metrics are detailed in the Experimental Setup Section). As shown in Figure 1, the resulting distribution across these levels is relatively balanced. This ensures that MedFact contains sufficient examples at varying levels of difficulty, enabling fine-grained differentiation of performance across a broad spectrum of model capabilities.
Model Zero-shot CoT
VC EL VC EL
F1 Precision Recall F1 F1 Precision Recall F1
Human 0.7521 0.7495 0.6588 0.7012
Open-source
II-Medical-8B 0.6361 0.3278 0.2618 0.2911 0.6477 0.2904 0.2240 0.2529
HuatuoGPT-o1-7B 0.5831 0.3054 0.1654 0.2146 0.5932 0.3508 0.1900 0.2465
DeepSeek-R1 0.6847 0.5035 0.7580 0.6051 0.6764 0.5012 0.7826 0.6111
DeepSeek-V3 0.6444 0.5832 0.3809 0.4608 0.6724 0.5266 0.5510 0.5386
Qwen2.5-72B-Instruct 0.5195 0.8024 0.3223 0.4599 0.6400 0.6731 0.4282 0.5234
Qwen3-235B-A22B 0.6944 0.6234 0.5803 0.6011 0.6468 0.6565 0.5076 0.5725
DeepSeek-R1-Distill-Llama-70B 0.6774 0.4990 0.7410 0.5964 0.6882 0.5142 0.7873 0.6221
DeepSeek-R1-Distill-Qwen-32B 0.5524 0.6648 0.3450 0.4543 0.5718 0.6480 0.3393 0.4454
Proprietary
XiaoYi 0.7126 0.6512 0.7023 0.6758 0.7061 0.6530 0.7221 0.6858
GPT-4o 0.6741 0.4966 0.5595 0.5262 0.6635 0.5012 0.6144 0.5520
GPT-4.5 0.6965 0.5694 0.6824 0.6208 0.6952 0.5971 0.6947 0.6422
o1 0.6713 0.6921 0.5652 0.6223 0.6810 0.6975 0.5775 0.6319
o3 0.6890 0.5355 0.8129 0.6456 0.6870 0.5579 0.8658 0.6785
Claude 3.7 Sonnet 0.6900 0.5242 0.5728 0.5474 0.6943 0.5567 0.6777 0.6113
Gemini 2.5 Flash 0.6787 0.4988 0.7921 0.6121 0.6730 0.4927 0.8299 0.6183
Gemini 2.5 Pro 0.6658 0.4689 0.8346 0.6005 0.6667 0.4828 0.8752 0.6223
Grok-4 0.6951 0.2675 0.2316 0.2482 0.6908 0.2136 0.1692 0.1888
Qwen-Max 0.7006 0.6113 0.5113 0.5569 0.6942 0.5663 0.6219 0.5928
ERNIE-X1 0.6792 0.5580 0.5775 0.5676 0.6871 0.6240 0.5945 0.6089
Doubao-Seed-1.6 0.7122 0.5566 0.6928 0.6173 0.7006 0.5609 0.7268 0.6332
Doubao-Seed-1.6-thinking 0.7139 0.6501 0.6938 0.6712 0.7050 0.6307 0.7344 0.6786
Table 1: Performance of different models on MedFact in zero-shot and CoT settings. For each metric, the best score is in bold and the second-best is underlined. Human expert performance is provided for reference.
Experiments
Experimental Setup
Models We evaluate a diverse set of 20 models on MedFact, comprising 8 open-source and 12 proprietary models:
  • Open-source: II-Medical-8B (Internet 2025), HuatuoGPT-o1-7B (Chen et al. 2024), DeepSeek-R1 (2025-05-28), DeepSeek-V3 (2025-03-24) (DeepSeek-AI 2025b), Qwen2.5-72B-Instruct (Team 2024), Qwen3-235B-A22B (Team 2025), DeepSeek-R1-Distill-Llama-70B (DeepSeek-AI 2025a), DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI 2025a).
  • Proprietary: XiaoYi (2025-06-24) (Healthcare 2025), GPT-4o (2024-11-20), GPT-4.5, o1, o3 (OpenAI 2025b), Claude 3.7 Sonnet, Gemini 2.5 Flash (DeepMind 2025), Gemini 2.5 Pro, Grok-4 (xAI 2025), Qwen2.5-Max (Team 2024), ERNIE-X1 (Baidu 2025), Doubao-Seed-1.6 (ByteDance 2025).
Among these models, XiaoYi, HuatuoGPT-o1-7B, and II-Medical-8B are specifically trained for the medical domain. We evaluate Doubao-Seed-1.6 in its standard version and thinking mode, listed as Doubao-Seed-1.6-thinking. We also establish a human expert baseline, calculated as the average performance of three medical professionals.
Evaluation Strategies We evaluate all models using zero-shot and Chain-of-Thought (CoT) prompting (Wei et al. 2022) in our main experiments. We also assess a subset of top-performing models using four advanced reasoning strategies, including recently proposed agent-based frameworks, which might be well-suited for fact-checking:
  • MedPrompt (Nori et al. 2023): A composite strategy that combines self-generated CoT, retrieval-augmented few-shot examples, and self-consistency (Wang et al. 2023b) for single-model inference.
  • Retrieval-Augmented Generation (RAG) (Lewis et al. 2020): A technique that grounds model outputs by first retrieving relevant documents from an external knowledge corpus to use as context in the prompt.
  • MAD (Liang et al. 2024): A multi-agent framework where agents engage in a multi-round debate to reach a consensus, thereby improving the final answer.
  • MDAgents (Kim et al. 2024): A dynamic medical framework that adaptively engages multi-agent collaboration based on the perceived difficulty of the input.
Evaluation Metrics We evaluate performance on two distinct tasks using Precision, Recall, and F1 score. The full implementation details are in the Appendix.
  • Veracity Classification (VC): This binary classification task evaluates a model’s ability to identify a medical text as either correct or incorrect. The “incorrect” label serves as the positive class in calculating metrics.
  • Error Localization (EL): For texts correctly classified as incorrect, this task evaluates the model’s ability to identify the span of text containing the error precisely. A prediction is considered a true positive only if the model localizes the correct error span.
We employ GPT-4o as the automated judge for the EL task. Its reliability is confirmed by strong agreement with human experts (Cohen’s Kappa = 0.87) on a random 10% data sample, validating its use as a proxy for human evaluation.
Model Veracity Classification Error Localization
Precision Recall F1 Precision Recall F1
DeepSeek-R1 0.5488 0.9101 0.6847 0.5035 0.7580 0.6051
   +MedPrompt 0.5657 0.9036 0.6958 0.5221 0.7580 0.6184
   +RAG (top-1) 0.6189 0.8658 0.7218 0.5892 0.7647 0.6656
   +RAG (top-3) 0.6393 0.8696 0.7369 0.6112 0.7713 0.6820
   +MAD 0.5310 0.9565 0.6829 0.4844 0.7940 0.6017
   +MDAgents 0.5411 0.9773 0.6965 0.4997 0.8280 0.6233
XiaoYi 0.6694 0.7618 0.7126 0.6512 0.7023 0.6758
   +MedPrompt 0.6899 0.7949 0.7387 0.6693 0.7231 0.6951
   +RAG (top-1) 0.7103 0.7788 0.7430 0.6900 0.7070 0.6984
   +RAG (top-3) 0.7179 0.7817 0.7484 0.6985 0.7117 0.7051
   +MAD 0.6088 0.8223 0.6996 0.6407 0.7316 0.6831
   +MDAgents 0.6038 0.8497 0.7059 0.5613 0.7136 0.6284
Table 2: Performance of DeepSeek-R1 and XiaoYi across different strategies on MedFact. For each model, the best score is in bold and the second-best is underlined.
Main Results
Overall Performance Table 1 summarizes the performance of all evaluated models. We report the F1 score for VC and Precision, Recall, and F1 for EL. Among the models evaluated, proprietary models consistently outperform their open-source counterparts across both tasks and all prompting strategies. Doubao-Seed-1.6-thinking emerges as the top-performing general-purpose model, while XiaoYi leads among medically specialized models. A significant performance gap remains between the best-performing models and human experts. The top model F1 score on the EL task, 0.6858, achieved by XiaoYi with CoT, is still below the human baseline of 0.7012. This disparity underscores both the difficulty of MedFact and the substantial challenge for current LLMs. The full results are provided in the Appendix.
A comparison between the tasks reveals a consistent performance gap, with all models performing worse on EL than on VC. For instance, the F1 scores for DeepSeek-V3 and Qwen2.5-Max decrease from 0.6947 and 0.7352 on VC to 0.5380 and 0.6116 on EL, respectively. This suggests that while models can often determine if a text contains an error, they struggle to pinpoint where it is. This difficulty often arises from a “correct-for-the-wrong-reason” phenomenon: a model may correctly classify a text as incorrect (a successful VC judgment) but do so by misidentifying a factually sound statement as the source of the error. Such behavior, rooted in a lack of deep medical knowledge, demonstrates that the EL task is a more stringent measure of a model’s fine-grained understanding and reveals fundamental deficiencies in its underlying medical knowledge.
Impact of Advanced Reasoning Strategies The results of DeepSeek-R1 and XiaoYi with advanced reasoning strategies against a zero-shot baseline are presented in Table 10. The retrieval corpus for RAG consists of the 6,405 expert-annotated source texts from our data curation process, excluding any rewrites. Results show that RAG using either one or three retrieved references yields substantial performance gains on both the VC and EL tasks. This finding shows that providing models with direct access to external, domain-specific knowledge is a highly effective method for medical fact-checking. MedPrompt delivers modest improvements, demonstrating that performance can also be enhanced solely through prompting techniques such as few-shot learning and self-consistency.
In contrast, the multi-agent frameworks (MAD and MDAgents) exhibit a consistent trade-off: they improve Recall at the expense of Precision, resulting in a stable F1 score for MAD and only a marginal improvement for MDAgents. We attribute this pattern to a phenomenon we term “over-criticism”. The collaborative process, which encourages the generation of multiple perspectives and error hypotheses, appears to increase the models’ propensity to flag errors where none exist, thereby elevating the false positive rate and lowering Precision. The slight performance edge of MDAgents may be attributed to its use of medical role-play, which could temper this critical tendency. This finding reveals a critical challenge: while these methods are designed to improve reasoning, they can inadvertently encourage an overly skeptical posture. This undermines reliability in medical domains, where correctly identifying accurate information is as crucial as flagging falsehoods.
Model F1 (VC) F1 (EL)
Qwen2.5-32B-Instruct 0.5990 0.4436
s1.1-32B 0.5941 0.3114
   + budget forcing 0.6002 0.3319
m1-32B-1K 0.6345 0.5082
   + budget forcing 0.6280 0.4926
Table 3: Performance of different models with and without inference-time scaling techniques on MedFact.
Analysis
Inference-time Scaling
Motivated by the superior performance of inference-time reasoning models such as DeepSeek-R1 and Doubao-Seed-1.6-thinking, over their models from the same series (DeepSeek-V3 and Doubao-Seed-1.6), we investigate the potential of inference-time scaling strategies. Specifically, we evaluate the impact of budget forcing (Muennighoff et al. 2025). This technique modulates a model’s reasoning process at inference time by either prematurely terminating generation or prolonging it via “Wait” tokens to adjust the computational budget. We focus on two models adapted for this paradigm: s1.1-32B and m1-32B-1K. s1.1-32B is a general reasoning model fine-tuned from Qwen2.5-32B-Instruct on only 1,000 high-quality reasoning examples (Muennighoff et al. 2025). m1-32B-1K is a specialized medical reasoning model fine-tuned from Qwen2.5-32B-Instruct, using 1,000 medical question-answering pairs (Huang et al. 2025).
Table 11 compares the performance of the baseline model (Qwen2.5-32B-Instruct) against s1.1-32B and m1-32B-1K under a CoT setting, evaluated both with and without budget forcing. The results indicate that s1.1-32B, fine-tuned on general-purpose data, suffers a substantial performance degradation on the EL task, which we attribute to its lack of domain-specific medical knowledge due to fine-tuning. In contrast, m1-32B-1K, with its medical fine-tuning, outperforms the base model on both tasks. However, applying budget forcing yields no performance gain and instead exacerbates the “over-criticism” phenomenon. The prolonged reasoning process causes models to hallucinate errors, a behavior analogous to the “over-thinking” observed in reasoning models (Ghosal et al. 2025; Sui et al. 2025). This suggests that simply extending a model’s reasoning time without providing new information can lead to counterproductive deliberation rather than improved accuracy. Ultimately, this finding reinforces the conclusion that for knowledge-intensive tasks like medical fact-checking, enhancing a model’s underlying medical knowledge is more critical than simply increasing its computational budget at inference.
Figure 4: Error distribution of DeepSeek-R1 and XiaoYi on the EL task (zero-shot).
Error Analysis
To better understand the performance gap between LLMs and human experts, and to diagnose why the EL task is particularly challenging, our human experts manually annotated and categorized the errors made by the top-performing models. As illustrated in Figure 4, the lack of medical knowledge constitutes the predominant category of model mistakes (76.09% of all errors for DeepSeek-R1), where models demonstrate insufficient understanding of medical principles, conflating distinct conditions or treatments, and relying on outdated information, which indicates that models often struggle with the nuanced, specialized knowledge required for accurate medical fact-checking.
Fine-grained Performance
We conduct a fine-grained analysis of model performance across different writing styles (Figure 9). For the VC task, models are more effective at identifying errors in fabricated medical misinformation than in texts from encyclopedias. This indicates that models can readily detect exaggerated falsehoods but struggle with the subtle, nuanced inaccuracies often present in professionally curated content. Notably, this performance gap diminishes for the more demanding EL task. This shift suggests that performance on localization is less influenced by stylistic cues and more fundamentally dependent on the model’s underlying medical knowledge. Consequently, the EL task serves as a more direct and robust measure of a model’s actual domain expertise.
Model R-2 F1 R-L F1
DeepSeek-R1 0.0300 0.0859
Qwen3-235B-A22B 0.0471 0.1201
XiaoYi 0.0450 0.1157
o3 0.0270 0.0767
Doubao-Seed-1.6-thinking 0.0265 0.0761
Table 4: ROUGE scores of different models on the text completion task.
Figure 5: Zero-shot performance of different models on MedFact across different writing styles.
Data Contamination
To investigate potential data contamination within the pre-training corpora of the evaluated LLMs, we devised a text completion task. For each instance in MedFact, we prompted the models with the first half of the text and tasked them with generating the continuation. We then measured the textual overlap between the generated output and the ground-truth second half using ROUGE-F1 scores, where a high score would suggest memorization. As shown in Table 12, all evaluated models, even those that achieve state-of-the-art performance on our tasks, exhibit negligible ROUGE scores. This result strongly indicates that MedFact serves as a valid and previously unseen benchmark for these models.
Conclusion
We introduce MedFact, a new benchmark for evaluating the fact-checking abilities of LLMs in Chinese medical texts. Constructed via a meticulous hybrid AI-human pipeline, our dataset is diverse, realistic, and uncontaminated, allowing it to probe for nuanced medical understanding. Our comprehensive evaluation of 20 leading LLMs reveals a critical performance deficit compared to human experts, particularly on the fine-grained task of error localization. Furthermore, LLMs exhibit an “over-criticism” failure mode that significantly degrades their precision. The challenges highlighted underscore the substantial work required to develop LLMs that are not just knowledgeable but also factually reliable enough for safe deployment in medical applications.
Appendix A Ethics Statement
Human Annotation and Compensation All expert annotators who contributed to MedFact received fair compensation for their time and expertise at rates that meet or exceed the minimum wage in their respective regions.
Privacy and Consent We employed a rigorous anonymization process to remove all Personally Identifiable Information (PII) and ensure privacy protection. Our use of this data is strictly for non-commercial research purposes and complies with the terms of service of source platforms under fair use principles.
Risk of Misinformation The factually incorrect instances in MedFact could be exploited to train models that generate harmful content or to disseminate misinformation directly. We will distribute the dataset under a restrictive license that explicitly prohibits use in clinical or patient-facing applications. The dataset will be accompanied by comprehensive documentation that clearly warns users about the presence of factual errors and specifies that its intended purpose is exclusively for research and benchmarking applications.
Bias and Representativeness MedFact focuses exclusively on Chinese medical texts. Consequently, the data, error distributions, and model findings may not generalize to other languages, cultures, or healthcare systems. We caution researchers against drawing overly broad conclusions from our findings and encourage the development of analogous benchmarks for other languages and cultural contexts.
Appendix B Dataset Construction & Statistics
Multi-criteria Filtering
Across the three filtering rounds, the acceptance rates (ratio of passed instances to candidate pool size) were 67.69%, 37.00%, and 23.62%, respectively. The consistency between the model suite’s judgment and medical professionals increased progressively: 69.72%, 84.37%, and 96.40%, respectively, while the reference database size grew from 412 to 624 instances. Figure 10 presents one accepted example, and Figure 11 presents three rejected examples from the reference database during the multi-criteria filtering process, illustrating the filtering criteria: overly simplistic, esoteric, and malformed content.
Error Taxonomy Definitions
We define eight fine-grained error types in MedFact, organized into two primary categories: Medical Errors and Non-Medical Errors. The definitions for each subcategory are detailed below.
Model Full Name
II-Medical-8B Intelligent-Internet/II-Medical-8B
HuatuoGPT-o1-7B FreedomIntelligence/HuatuoGPT-o1-7B
DeepSeek-R1 deepseek-ai/DeepSeek-R1-0528
DeepSeek-V3 deepseek-ai/DeepSeek-V3-0324
Qwen2.5-72B-Instruct Qwen/Qwen2.5-72B-Instruct
Qwen3-235B-A22B Qwen/Qwen3-235B-A22B
DeepSeek-R1-Distill-Llama-70B deepseek-ai/DeepSeek-R1-Distill-Llama-70B
DeepSeek-R1-Distill-Qwen-32B deepseek-ai/DeepSeek-R1-Distill-Qwen-32B
XiaoYi xiaoyi-2025-06-24
GPT-4o gpt-4o-2024-11-20
GPT-4.5 gpt-4.5-preview-2025-02-27
o1 o1-2024-12-17
o3 o3-2025-04-16
Claude 3.7 Sonnet claude-3-7-sonnet-20250219
Gemini 2.5 Flash gemini-2.5-flash
Gemini 2.5 Pro gemini-2.5-pro
Grok-4 grok-4-0709
Qwen2.5-Max qwen-max-2025-01-25
ERNIE-X1 ernie-x1-32k-preview
Doubao-Seed-1.6 doubao-seed-1.6
Doubao-Seed-1.6-thinking doubao-seed-1.6-thinking
Table 5: Model versions and full names of the evaluated models.
Medical Errors Medical errors encompass factually incorrect content related to established medical knowledge, principles, or clinical practice. The definitions of each error type in MedFact are as follows:
  • Conceptual Error: Incorrect explanations of medical concepts, disease mechanisms, or drug actions; treatment recommendations that contradict established clinical guidelines; or inaccurate medication instructions (e.g., dosage, frequency). This category also encompasses errors in standard reference values for clinical data, such as laboratory test ranges or clinical scoring systems (e.g., GCS, APGAR).
  • Terminological Error: Use of non-standard, obsolete, or erroneous medical terminology. This includes incorrect translations or mappings between medical terms across different languages (e.g., Chinese and English).
  • Temporal Error: Errors in the chronological sequence of medical events or processes. Examples include misordering clinical procedures (e.g., administering post-operative medication pre-operatively), misrepresenting disease progression timelines, or describing developmental stages out of sequence.
  • Others (Medical): A residual category for medical errors not classifiable within the above types. This includes failures to convey critical safety warnings, flawed clinical reasoning (e.g., broken causal chains between symptoms and diagnoses), or misrepresentation of risk factors.
Non-Medical Errors Non-medical errors include factual inaccuracies that are not domain-specific to medicine.
  • Citation and Sourcing Error: Fabrication or misrepresentation of sources, including citing non-existent literature, providing inaccurate citation details, or inventing data sources.
  • Biased Content Error: Content that reflects or promotes bias by perpetuating stereotypes related to social identity in diagnoses or treatment recommendations, or employing language that reinforces social prejudice.
  • General Factual Error: Factual inaccuracies in non-medical background information, encompassing errors related to history, geography, basic scientific principles, or general world knowledge.
  • Others (Non-medical): A residual category for other non-medical issues, including ethical or legal violations (e.g., privacy breaches, suggesting illegal acts), undisclosed commercial promotion, or severe linguistic incoherence that fundamentally obstructs comprehension.
Writing Style Taxonomy Definitions
We categorize the writing styles of MedFact into four distinct types: encyclopedia-style articles, popular science journalism, user-generated web content, and fabricated misinformation:
  • Encyclopedia-style: Formal content characteristic of medical texts, featuring specialized terminology and structured presentation.
  • Popular Science Journalism: Articles that translate complex medical topics for lay audiences, employing simplified language, narrative elements, and references to purported experts to establish credibility.
  • User-generated Web Content: Informal medical discussions sourced from online Q&A platforms, characterized by colloquial language and personal anecdotes.
  • Fabricated Misinformation: Deliberately false or misleading medical content that employs exaggerated claims and distorts facts, featuring sensationalized language, cherry-picked data, and conspiracy theories while mimicking legitimate sources.
Error Category Count Proportion (%)
I. Medical Errors
   Conceptual 557 52.65
   Terminological 236 22.31
   Temporal 119 11.25
   Other Medical 34 3.21
II. Non-Medical Errors
   Citation & Sourcing 38 3.59
   Biased Content 37 3.50
   General Factual 24 2.27
   Other Non-Medical 13 1.23
Table 6: Distribution of 8 fine-grained error types among the 1,058 incorrect instances in MedFact.
Accuracy Interval Count Proportion (%)
0–20% 219 10.35
20–40% 298 14.08
40–60% 468 22.12
60–80% 600 28.36
80–100% 531 25.09
Table 7: Distribution of samples in MedFact, categorized by accuracy score intervals.
Data Statistics
Tables 6 and 7 present the data distribution across different error types and difficulty levels, respectively. Difficulty levels are determined based on model performance.
Data Examples
Figures 14 to 17 present examples for each medical error. Figures 18 to 21 present examples for each non-medical error. Figures 22 to 25 present examples for each writing style.
Appendix C Implementation Details
Prompt Templates
Figures 12 to 13 present the prompts used for zero-shot and CoT prompting. All prompts were originally formulated in Chinese; we provide their corresponding English translations for clarity and reference.
Model Versions
We accessed proprietary models through API calls and deployed open-source models locally. Table 5 lists the model versions and full names used in our experiments.
LLM-as-a-Judge
For the evaluation of error localization tasks, we assessed the consistency agreement between different models and human experts. DeepSeek-R1, GPT-4o, o1, Claude 3.7 Sonnet, and Gemini 2.5 Pro achieved Cohen’s Kappa values of 0.800, 0.870, 0.870, 0.845, and 0.865, respectively, on a random 10% sample of the veracity classification task results from Doubao-Seed-1.6-thinking, XiaoYi, and Qwen-Max. Considering API costs and the evaluation performance, we selected GPT-4o for the LLM-as-a-judge evaluation.
Appendix D Complete Experimental Results
Full Result Tables
We present the complete results of our main experiments in Tables 8 and 9 and Figures 6 and 7. The results from experiments using advanced evaluation strategies and inference-time scaling are provided in Tables 10 and 11, respectively.
Model Zero-shot CoT
Precision Recall F1 Score Precision Recall F1 Score
Human 0.7695 0.7353 0.7521
Open-source
II-Medical-8B 0.5687 0.7216 0.6361 0.5724 0.7459 0.6477
HuatuoGPT-o1-7B 0.6004 0.5668 0.5831 0.6185 0.5699 0.5932
Deepseek-R1 0.5488 0.9101 0.6847 0.5360 0.9163 0.6764
Deepseek-V3 0.6883 0.6057 0.6444 0.6006 0.7636 0.6724
Qwen2.5-72B-Instruct 0.8264 0.3788 0.5195 0.7320 0.5686 0.6400
Qwen3-235B-A22B 0.6720 0.7183 0.6944 0.6949 0.6049 0.6468
DeepSeek-R1-Distill-Llama-70B 0.5451 0.8947 0.6774 0.5490 0.9220 0.6882
DeepSeek-R1-Distill-Qwen-32B 0.7204 0.4480 0.5524 0.7194 0.4744 0.5718
Proprietary
XiaoYi 0.6694 0.7618 0.7126 0.6631 0.7552 0.7061
GPT-4o 0.5842 0.7968 0.6741 0.5646 0.8044 0.6635
GPT-4.5 0.6108 0.8100 0.6965 0.6254 0.7826 0.6952
o1 0.7155 0.6323 0.6713 0.7205 0.6456 0.6810
o3 0.5596 0.8960 0.6890 0.5507 0.9128 0.6870
Claude 3.7 Sonnet 0.6063 0.8006 0.6900 0.6021 0.8197 0.6943
Gemini 2.5 Flash 0.5369 0.9225 0.6787 0.5240 0.9405 0.6730
Gemini 2.5 Pro 0.5067 0.9707 0.6658 0.5082 0.9688 0.6667
Grok-4 0.5785 0.8705 0.6951 0.5789 0.8563 0.6908
Qwen-Max 0.6873 0.7146 0.7006 0.6031 0.8175 0.6942
ERNIE-X1 0.6210 0.7495 0.6792 0.6643 0.7116 0.6871
Doubao-Seed-1.6 0.6086 0.8582 0.7122 0.5979 0.8459 0.7006
Doubao-Seed-1.6-thinking 0.6711 0.7625 0.7139 0.6437 0.7791 0.7050
Table 8: Veracity classification performance of human experts and models on MedFact in zero-shot and CoT settings. The best results are bolded and the second-best are underlined.
Figure 6: Veracity classification performance of human experts and models on MedFact in zero-shot and CoT settings.
Model Zero-shot CoT
Precision Recall F1 Score Precision Recall F1 Score
Human 0.7495 0.6588 0.7012
Open-source
II-Medical-8B 0.3278 0.2618 0.2911 0.2904 0.2240 0.2529
HuatuoGPT-o1-7B 0.3054 0.1654 0.2146 0.3508 0.1900 0.2465
Deepseek-R1 0.5035 0.7580 0.6051 0.5012 0.7826 0.6111
Deepseek-V3 0.5832 0.3809 0.4608 0.5266 0.5510 0.5386
Qwen2.5-72B-Instruct 0.8024 0.3223 0.4599 0.6731 0.4282 0.5234
Qwen3-235B-A22B 0.6234 0.5803 0.6011 0.6565 0.5076 0.5725
DeepSeek-R1-Distill-Llama-70B 0.4990 0.7410 0.5964 0.5142 0.7873 0.6221
DeepSeek-R1-Distill-Qwen-32B 0.6648 0.3450 0.4543 0.6480 0.3393 0.4454
Proprietary
XiaoYi 0.6512 0.7023 0.6758 0.6530 0.7221 0.6858
GPT-4o 0.4966 0.5595 0.5262 0.5012 0.6144 0.5520
GPT-4.5 0.5694 0.6824 0.6208 0.5971 0.6947 0.6422
o1 0.6921 0.5652 0.6223 0.6975 0.5775 0.6319
o3 0.5355 0.8129 0.6456 0.5579 0.8658 0.6785
Claude 3.7 Sonnet 0.5242 0.5728 0.5474 0.5567 0.6777 0.6113
Gemini 2.5 Flash 0.4988 0.7921 0.6121 0.4927 0.8299 0.6183
Gemini 2.5 Pro 0.4689 0.8346 0.6005 0.4828 0.8752 0.6223
Grok-4 0.2675 0.2316 0.2482 0.2136 0.1692 0.1888
Qwen-Max 0.6113 0.5113 0.5569 0.5663 0.6219 0.5928
ERNIE-X1 0.5580 0.5775 0.5676 0.6240 0.5945 0.6089
Doubao-Seed-1.6 0.5566 0.6928 0.6173 0.5609 0.7268 0.6332
Doubao-Seed-1.6-thinking 0.6501 0.6938 0.6712 0.6307 0.7344 0.6786
Table 9: Error localization performance of human experts and models on MedFact in zero-shot and CoT settings. The best results are bolded and the second-best are underlined.
Figure 7: Error localization performance of human experts and models on MedFact in zero-shot and CoT settings.
Model Veracity Classification Error Localization
Precision Recall F1 Score Precision Recall F1 Score
Open-source
DeepSeek-R1 0.5488 0.9101 0.6847 0.5035 0.7580 0.6051
   +MedPrompt 0.5657 0.9036 0.6958 0.5221 0.7580 0.6184
   +RAG (top-1) 0.6189 0.8658 0.7218 0.5892 0.7647 0.6656
   +RAG (top-3) 0.6393 0.8696 0.7369 0.6112 0.7713 0.6820
   +MAD 0.5310 0.9565 0.6829 0.4844 0.7940 0.6017
   +MDAgents 0.5411 0.9773 0.6965 0.4997 0.8280 0.6233
Qwen3-235B-A22B 0.6720 0.7183 0.6944 0.6234 0.5803 0.6011
   +MedPrompt 0.6540 0.7344 0.6919 0.6233 0.6427 0.6329
   +RAG (top-1) 0.8380 0.5964 0.6969 0.8222 0.5331 0.6468
   +RAG (top-3) 0.7646 0.6938 0.7275 0.7466 0.6295 0.6831
   +MAD 0.6695 0.7335 0.7000 0.6219 0.5955 0.6084
   +MDAgents 0.6628 0.7637 0.7097 0.6137 0.6172 0.6155
Proprietary
XiaoYi 0.6694 0.7618 0.7126 0.6512 0.7023 0.6758
   +MedPrompt 0.6899 0.7949 0.7387 0.6693 0.7231 0.6951
   +RAG (top-1) 0.7103 0.7788 0.7430 0.6900 0.7070 0.6984
   +RAG (top-3) 0.7179 0.7817 0.7484 0.6985 0.7117 0.7051
   +MAD 0.6088 0.8223 0.6996 0.6407 0.7316 0.6831
   +MDAgents 0.6038 0.8497 0.7059 0.5613 0.7136 0.6284
o3 0.5596 0.8960 0.6890 0.5355 0.8129 0.6456
   +MedPrompt 0.6357 0.7883 0.7038 0.6126 0.7146 0.6597
   +RAG (top-1) 0.6152 0.8658 0.7193 0.6013 0.8166 0.6926
   +RAG (top-3) 0.6202 0.8658 0.7227 0.6024 0.8034 0.6885
   +MAD 0.5495 0.9178 0.6874 0.5221 0.8611 0.6500
   +MDAgents 0.5556 0.9216 0.6932 0.5376 0.8573 0.6608
Doubao-Seed-1.6-thinking 0.6711 0.7625 0.7139 0.6501 0.6938 0.6712
   +MedPrompt 0.6786 0.7543 0.7144 0.6628 0.7023 0.6820
   +RAG (top-1) 0.7680 0.7164 0.7413 0.7595 0.6834 0.7194
   +RAG (top-3) 0.7695 0.7178 0.7428 0.7605 0.6815 0.7188
   +MAD 0.6352 0.8129 0.7131 0.6132 0.7401 0.6707
   +MDAgents 0.6372 0.8516 0.7290 0.6261 0.8119 0.7070
Table 10: Performance of different models across different strategies on MedFact. For each model, the best results are bolded and second-best are underlined.
Figure 8: Changes in precision and recall after applying the MAD and MDAgents frameworks.
Model Veracity Classification Error Localization
Precision Recall F1 Score Precision Recall F1 Score
Qwen2.5-32B-Instruct 0.7192 0.5132 0.5990 0.6307 0.3422 0.4436
s1.1-32B 0.6378 0.5560 0.5941 0.4369 0.2420 0.3114
   + budget forcing 0.5120 0.7250 0.6002 0.3275 0.3365 0.3319
m1-32B-1K 0.6540 0.6163 0.6345 0.5808 0.4518 0.5082
   + budget forcing 0.5953 0.6645 0.6280 0.5122 0.4745 0.4926
Table 11: Performance of different models with and without test-time scaling (budget forcing) on MedFact. The best results are bolded and the second-best are underlined.
Model ROUGE-1 F1 ROUGE-2 F1 ROUGE-L F1 LCS Normalized Avg
DeepSeek-R1 0.0923 0.0300 0.0859 0.0458
Qwen3-235B-A22B 0.1286 0.0471 0.1201 0.0456
XiaoYi 0.1236 0.0450 0.1157 0.0448
GPT-4.5 0.1548 0.0549 0.1435 0.0401
o1 0.1449 0.0548 0.1365 0.0442
o3 0.0817 0.0270 0.0767 0.0437
Gemini 2.5 Pro 0.1267 0.0440 0.1176 0.0490
Doubao-Seed-1.6 0.1327 0.0495 0.1246 0.0465
Doubao-Seed-1.6-thinking 0.0819 0.0265 0.0761 0.0455
Table 12: ROUGE scores of different models on the text completion task.
Fine-grained Performance
Figure 9 details the fine-grained model performance across various error types, medical specialties, and writing styles.
Figure 9: Performance breakdown of different models on MedFact in the zero-shot setting. The radar charts illustrate model performance across three key dimensions: error type, medical specialty, and textual style.
Data Contamination
Table 12 presents the results of our data contamination analysis for the task completion experiment.
Appendix E Case Studies
We categorize the observed errors using the following typology:
  • Conceptual Errors: Fundamental misunderstanding of medical principles, such as incorrectly identifying a disease’s etiology or misinterpreting mechanisms.
  • Concept Conflation: Conflation of distinct but related medical concepts, such as mistaking the symptoms of one disease for those of a similar-presenting condition, or confusing two drugs with similar names but different therapeutic effects.
  • Outdated Information: Reliance on information that is no longer current, including citation of superseded clinical guidelines, obsolete treatment protocols, or outdated pharmacological data.
  • Others (Medical): Domain-specific errors not captured by the preceding categories, such as dosage miscalculations, inappropriate procedural indications, or misinterpretation of diagnostic criteria.
  • Failure to Follow Instructions: Non-adherence to prompt specifications, such as generating output in a format inconsistent with the given instructions.
  • Logical Errors: Errors in reasoning where a model’s explanation for its judgment is internally inconsistent or contains logical contradictions.
  • Others (Non-medical): General errors unrelated to medical expertise, such as grammatical mistakes, factual inaccuracies in non-medical information, or other fundamental reasoning flaws.
Figure 26 and  27 present successful fact-checking cases. Figures 28 to 41 present case studies for each error category. We further illustrate the “correct-for-the-wrong-reason” and “over-criticism” phenomena with case studies in Figures 42 to 46.
Appendix F Limitations & Future Work
Linguistic and Cultural Scope Our benchmark is developed exclusively for the Chinese language and is situated within the context of the Chinese healthcare system and common public discourse. The distribution and performance of LLMs may not generalize directly to other languages, cultures, or healthcare systems. This focus represents a deliberate methodological choice to ensure high linguistic quality and cultural consistency. Future work could involve creating parallel benchmarks in other languages to enable cross-lingual comparative studies.
Single Error Assumption For clarity of annotation and evaluation, each incorrect text in MedFact contains only a single factuality error. This represents a simplifying assumption, as real-world misinformation often contains multiple, interwoven inaccuracies. Extending the benchmark to include instances with multiple, interacting errors would provide a more realistic and challenging evaluation of LLM fact-checking capabilities.
Reliance on LLM-as-a-Judge For the EL task, we employ GPT-4o as an automated judge. While this enables scalable evaluation, it is subject to the inherent limitations of LLM judges, including potential biases and inconsistencies. Future research could explore hybrid evaluation frameworks that combine the scalability of LLM judges with targeted, expert-led validation for ambiguous or high-stakes cases.
Static Knowledge MedFact represents a snapshot of medical knowledge at the time of its creation. As medical science evolves, some information may become outdated. Although we have prioritized foundational medical knowledge to minimize this issue, we acknowledge that establishing a process for dynamically updating the benchmark would constitute a valuable long-term objective.
Figure 10: An example of accepted data with corresponding professional feedback.
Figure 11: Examples of rejected data with corresponding professional feedback.
Figure 12: Prompt template used for zero-shot prompting.
Figure 13: Prompt template used for CoT prompting.
Figure 14: A data example of an error categorized as medical conceptual error.
Figure 15: A data example of an error categorized as medical terminological error.
Figure 16: A data example of an error categorized as medical temporal error.
Figure 17: A data example of an error categorized as other medical error.
Figure 18: A data example of an error categorized as citation & sourcing error.
Figure 19: A data example of an error categorized as biased content.
Figure 20: A data example of an error categorized as general factual error.
Figure 21: A data example of an error categorized as other non-medical error.
Figure 22: A data example of an error categorized as encyclopedia-style.
Figure 23: Data examples with writing style of popular science journalism.
Figure 24: Data examples with writing style of user-generated web content.
Figure 25: Data examples with writing style of fabricated misinformation.
Figure 26: Case of successful detection of factual correctness.
Figure 27: Case of successful detection of factual correctness.
Figure 28: Failure case with conceptual errors.
Figure 29: Failure case with conceptual errors.
Figure 30: Failure case with concept conflation.
Figure 31: Failure case with concept conflation.
Figure 32: Failure case with outdated information.
Figure 33: Failure case with outdated information.
Figure 34: Failure case with other medical errors.
Figure 35: Failure case with other medical errors.
Figure 36: Failure case with failure to follow instruction errors.
Figure 37: Failure case with failure to follow instruction errors.
Figure 38: Failure case with logical errors.
Figure 39: Failure case with logical errors.
Figure 40: Failure case with other non-medical errors.
Figure 41: Failure case with other non-medical errors.
Figure 42: Failure case with correct-for-the-wrong-reason errors.
Figure 43: Failure case with over-criticism errors.
Figure 44: Failure case with over-criticism errors.
Figure 45: Failure case with over-criticism errors under the budget forcing technique.
Figure 46: Failure case with over-criticism errors under the MAD framework.

### v2

@@
 Abstract
-The increasing deployment of Large Language Models (LLMs) in healthcare necessitates a rigorous evaluation of their factual reliability.
-However, existing benchmarks are often limited by narrow domains of data, failing to capture the complexity of real-world medical information.
-To address this critical gap, we introduce MedFact, a new and challenging benchmark for Chinese medical fact-checking.
-MedFact comprises 2,116 expert-annotated instances curated from diverse real-world texts, spanning 13 medical specialties, 8 fine-grained error types, 4 writing styles, and multiple difficulty levels.
-Its construction employs a hybrid AI-human framework where iterative expert feedback refines an AI-driven, multi-criteria filtering process, ensuring both high data quality and difficulty.
-We conduct a comprehensive evaluation of 20 leading LLMs, benchmarking their performance on veracity classification and error localization against a human expert baseline.
-Our results reveal that while models can often determine if a text contains an error, precisely localizing it remains a substantial challenge, with even top-performing models falling short of human performance.
-Furthermore, our analysis uncovers a frequent “over-criticism” phenomenon, a tendency for models to misidentify correct information as erroneous, which is exacerbated by advanced reasoning techniques such as multi-agent collaboration and inference-time scaling.
-By highlighting these critical challenges for deploying LLMs in medical applications, MedFact provides a robust resource to drive the development of more factually reliable and medically aware models.
-Content warning: This paper contains content that may be inappropriate or offensive.
+Deploying Large Language Models (LLMs) in medical applications requires fact-checking capabilities to ensure patient safety and regulatory compliance.
+We introduce MedFact, a challenging Chinese medical fact-checking benchmark with 2,116 expert-annotated instances from diverse real-world texts, spanning 13 specialties, 8 error types, 4 writing styles, and 5 difficulty levels.
+Construction uses a hybrid AI-human framework where iterative expert feedback refines AI-driven, multi-criteria filtering to ensure high quality and difficulty.
+We evaluate 20 leading LLMs on veracity classification and error localization, and results show models often determine if text contains errors but struggle to localize them precisely, with top performers falling short of human performance.
+Our analysis reveals the “over-criticism” phenomenon, a tendency for models to misidentify correct information as erroneous, which can be exacerbated by advanced reasoning techniques such as multi-agent collaboration and inference-time scaling.
+MedFact highlights the challenges of deploying medical LLMs and provides resources to develop factually reliable medical AI systems.
+Warning: This paper contains content that may be inappropriate or offensive.
+MedFact: Benchmarking the Fact-Checking
+Capabilities of Large Language Models on Chinese Medical Texts
+Jiayi He, Yangmin Huang, Qianyun Du, Xiangying Zhou, Zhiyang He, Jiaxue Hu, Xiaodong Tao, Lixian Lai Xunfei Healthcare Technology Co., Ltd.
+{jyhe14, ymhuang9, qydu, xyzhou30, zyhe, jxhu2, xdtao, lxlai2}@iflytek.com
 footnotetext: Corresponding authors.
-Our data and code: https://github.com/ivy3h/MedFact.
-Introduction
-Large Language Models (LLMs) (OpenAI 2024a; DeepSeek-AI 2025a; He et al.
-2025; Tang et al.
-2025), renowned for their versatile capabilities in natural language processing (NLP), are increasingly being applied to real-world medical applications (Huang et al.
-2024; Wu et al.
-2025), such as clinical decision support (Rao et al.
-2023), patient assessment (Zack et al.
-2024), diagnosis (Kuroiwa et al.
-2023), and text classification (Zhang et al.
-2022).
-While various benchmarks exist to assess the medical capabilities of LLMs, they typically focus on tasks like medical question answering (QA)(Jin et al.
-2021) and relation extraction(Guan et al.
-2020).
-However, the crucial issue of their factual reliability in the medical domain remains underexplored.
-This gap is particularly alarming given the potential for severe harm from medical misinformation (Wang and Togher 2024).
-Existing efforts to bridge this gap are limited.
-For instance, VeriFact (Chung et al.
-2025) evaluates fact-checking on synthetic, LLM-generated clinical text.
-MEDEC (Abacha et al.
-2025) is restricted to error detection in clinical notes.
-By focusing on synthetic data or a single text genre, these benchmarks fail to capture the diversity of medical information encountered in the real world.
-This narrow focus precludes a comprehensive assessment of LLM fact-checking capabilities and highlights the need for a benchmark covering a broader range of realistic medical content.
-To address these limitations, we introduce MedFact, a challenging benchmark for Chinese medical fact-checking, which is built upon three core principles: rigorous design, broad coverage, and uncontaminated evaluation.
+Project Page.
+1 Introduction
+Large Language Models (LLMs) openai2024gpt4technicalreport; deepseekai2025deepseekr1incentivizingreasoningcapability; he-etal-2025-self; 11030757, with their strong natural language processing capabilities, are increasingly applied to medical tasks huang2024comprehensivesurveyevaluatinglarge; wu-etal-2025-believe, such as clinical decision support RAO2023990 and patient assessment zack2024assessing.
+Medical texts sourced from the internet are also widely integrated into Retrieval-Augmented Generation (RAG) frameworks for medical dialog AI systems in real-world deployments.
+However, these texts often contain factual errors, and their use in clinical consultation scenarios poses severe risks, including patient harm and legal liability.
+Ensuring the factual accuracy of medical content is essential for the responsible deployment of LLM-based medical systems 10.1001/jamapediatrics.2023.5282.
+Despite this importance, existing benchmarks primarily evaluate LLM performance on tasks such as medical question answering (QA) app11146421 and relation extraction 10.1007/978-3-030-60450-9_22, leaving their factual reliability largely underexplored.
+Current efforts to address this gap remain limited.
+For instance, VeriFact Chung2025 evaluates fact-checking on synthetic, LLM-generated clinical text, while MEDEC abacha2025medecbenchmarkmedicalerror is restricted to error detection in clinical notes.
+By focusing on synthetic data or a single text genre, these benchmarks fail to capture the diversity of medical information encountered in real-world medical AI systems.
+This limitation hinders the development of trustworthy medical AI systems, underscoring the need for a benchmark that encompasses a broader range of realistic medical content to facilitate a comprehensive evaluation of LLM fact-checking capabilities.
+Figure 1: An overview of the data construction pipeline and the components of MedFact.
+To address these issues, we introduce MedFact, a challenging benchmark for Chinese medical fact-checking, built upon three core principles:
   • Rigorously Designed Pipeline: We construct MedFact through a human-in-the-loop pipeline that integrates large-scale AI filtering with fine-grained annotation by medical professionals.
@@
 It comprises 2,116 expert-annotated medical texts that span 13 medical specialties, 8 fine-grained error types, 4 writing styles, and multiple difficulty levels.
-  • Uncontaminated Data: MedFact is constructed from proprietary texts under copyright-compliant agreements, making it highly unlikely to be part of the common pre-training corpora for LLMs.
-This helps ensure a fairer assessment of their true fact-checking capabilities.
-We benchmark 20 leading proprietary and open-source LLMs on MedFact across two distinct tasks: veracity classification (VC) and the more demanding error localization (EL).
-Our experiments reveal that even state-of-the-art models exhibit significant limitations, with performance falling substantially short of human expert levels.
-Moreover, while retrieval-augmented generation (RAG) can enhance performance, we find that other strategies, such as multi-agent collaboration and inference-time budget forcing, induce an overly critical bias, causing models to misclassify correct texts as incorrect frequently.
-This finding underscores the critical gap between current LLM capabilities and the requirements for safe deployment in high-stakes medical applications.
+  • Uncontaminated Data: MedFact is constructed from proprietary texts, making it highly unlikely to appear in LLM pre-training data, which ensures a fairer assessment of their true fact-checking capabilities.
+We benchmark 20 LLMs on MedFact across veracity classification and error localization tasks.
+State-of-the-art models still perform below human experts.
+While RAG can improve performance, multi-agent collaboration and inference-time compute scaling induce overly critical bias, causing frequent misclassifications of correct texts as incorrect.
+This highlights the gap between current LLM capabilities and the requirements for safe medical deployment.
 Our main contributions are:
-  • We introduce MedFact, a novel benchmark for Chinese medical fact-checking, distinguished by its rigorous construction, broad coverage of real-world medical content, and its design to mitigate training data contamination.
-  • We present a comprehensive evaluation of 20 LLMs, revealing a significant performance gap relative to human experts on both veracity classification and the precise localization of factual errors.
-  • We conduct in-depth analyses of model behavior, examining the impact of advanced strategies like multi-agent collaboration and inference-time scaling.
-Our findings reveal an “over-criticism” phenomenon, and demonstrate that MedFact remains a challenge even for state-of-the-art models using sophisticated inference techniques.
-Figure 1: An overview of the data construction pipeline and the components of MedFact.
-Related Works
-LLMs in Medical Domains
-The significant capabilities of Large Language Models (LLMs) across many disciplines have prompted the exploration of their applications in the medical domain (Huang et al.
-2024; Chen et al.
-2025).
-Medprompt (Nori et al.
-2023) enhances LLMs’ performance on medical benchmarks by combining self-generated Chain-of-Thought (CoT) (Wei et al.
-2022), few-shot learning with dynamic example selection (Liu et al.
-2022), and self-consistency (Wang et al.
-2023b).
-Other developments include HuaTuo (Wang et al.
-2023a), a LLaMA-based model fine-tuned on Chinese medical instruction data; HuatuoGPT (Zhang et al.
-2023), a model specialized for medical consultation, fine-tuned on responses distilled from ChatGPT and real-world data; and PediatricsGPT (Yang et al.
-2024), a model developed via a systematic training pipeline for pediatric and general medical applications.
-These advances highlight a growing interest in adapting LLMs to the challenges of medical reasoning and the specialized requirements of real-world practice.
-Benchmarking Medical Knowledge in LLMs
-Numerous benchmarks have been developed to evaluate the medical knowledge of LLMs (Nguyen et al.
-2023; Kweon et al.
-2024; Abacha et al.
-2025; Wu et al.
-2025).
-MedQA (Jin et al.
-2021) is a multiple-choice QA dataset derived from the United States Medical Licensing Examinations to assess the clinical knowledge of LLMs.
-More recently, HealthBench (Arora et al.
-2025) is a benchmark of realistic health conversations evaluated with custom, physician-created rubrics.
-MedXpertQA (Zuo et al.
-2025) evaluates expert-level medical knowledge using both textual and multimodal tasks.
-While some benchmarks for medical fact-checking exist, they often focus on narrow contexts.
-VeriFact (Chung et al.
-2025) verifies synthetic text against structured Electronic Health Records using a retrieval-augmented generation (RAG) and LLM-as-a-Judge paradigm.
-MEDEC (Abacha et al.
-2025) provides a benchmark for error detection and correction, but is confined to clinical notes.
-Relying on synthetic data or a single text genre, these benchmarks do not capture the diverse contexts of medical information encountered in the real world, limiting their utility for assessing how well models generalize.
-Consequently, a research gap remains for a benchmark that can robustly evaluate fact-checking across a wide array of realistic medical texts.
-Benchmarking Factuality in LLMs
-Recent research has produced several benchmarks to evaluate factuality in LLMs.
-For instance, HaluEval (Li et al.
-2023) focuses on hallucination detection through a large-scale collection of automatically generated and human-annotated examples.
-SimpleQA (Wei et al.
-2024) measures short-form factuality using single-answer questions adversarially collected to challenge state-of-the-art models like GPT-4.
-While these benchmarks provide valuable tools for assessing general-domain factuality, they lack the domain-specific nuances required for high-stakes fields like medicine, where misinformation can have severe public health consequences (Wang and Togher 2024).
-Addressing this challenge requires a solution that bridges two critical gaps: the lack of domain-specific context in general factuality benchmarks and the lack of real-world diversity in existing medical fact-checking benchmarks.
-To this end, we introduce MedFact, a new benchmark explicitly designed to evaluate the fact-checking capabilities of LLMs on a comprehensive and diverse collection of realistic medical texts.
-MedFact
-Data Collection
-The construction pipeline for MedFact is illustrated in Figure 1.
-We began with a corpus of 27,116 medical texts secured through a copyright-compliant data-sharing agreement with a commercial partner.
-The source materials are primarily drawn from the partner’s internal medical encyclopedia, supplemented with content from its medical consultant platforms, including question-answering pages and user forum discussions.
-As these sources have not been previously released to the public, this approach substantially mitigates the risk of training data contamination.
-Data Curation
-Multi-criteria Filtering We defined a suite of seven leading LLMs: GPT-4o (OpenAI 2024b), GPT-4.5 (OpenAI 2025a), o1 (OpenAI 2024c), Gemini 2.5 Pro (DeepMind 2025), DeepSeek-R1 (DeepSeek-AI 2025a), Claude 3.7 Sonnet (Anthropic 2025), and Doubao-Seed-1.6-thinking (ByteDance 2025).
-This model suite formed an ensemble that, via majority vote, filtered the initial corpus by discarding texts identified as overly simplistic (e.g., lacking medical knowledge), esoteric (e.g., highly specialized content), or malformed (e.g., missing a clear subject or context).
-This automated process was refined through a three-round, iterative human-in-the-loop feedback mechanism inspired by Wu et al.
-(2024).
-In each round, medical experts reviewed a 5% sample of filtered data.
-We then used their annotations on misclassified items to construct a more effective retrieval-augmented, few-shot prompt for the subsequent round.
-This prompt leveraged criterion-specific retrieval, augmenting each input with the most semantically similar expert-annotated example for each of our three filtering criteria: simplistic, esoteric, and malformed.
-Over three rounds, the data acceptance rate decreased from 67.69% to 37.00%, and finally to 23.62%.
-The resulting filter achieved 96.40% agreement with expert judgments, reducing the candidate pool from 27,116 to 6,405 texts.
+  • We introduce MedFact, a novel Chinese medical fact-checking benchmark with rigorous construction, broad real-world coverage, and a design that mitigates data contamination.
+  • We present an evaluation of 20 LLMs, revealing a significant performance gap relative to humans on veracity classification and the precise localization of factual errors.
+  • We analyze advanced strategies such as inference-time scaling and reveal a “over-criticism” phenomenon where models excessively flag correct information as erroneous.
+Figure 2: An example of fact-checking performed by Claude 3.7 Sonnet.
+Figure 3: Distribution of the specialties in MedFact.
+2 Related Works
+2.1 LLMs in Medical Domains
+LLMs have prompted extensive exploration of medical applications huang2024comprehensivesurveyevaluatinglarge.
+Medprompt nori2023generalistfoundationmodelsoutcompete enhances medical performance by combining self-generated Chain-of-Thought (CoT)NEURIPS2022_9d560961, few-shot learning with dynamic examplesliu-etal-2022-makes, and self-consistency wangself.
+Other developments include HuaTuo wang2023huatuotuningllamamodel, a LLaMA-based model fine-tuned on Chinese medical instructions, and PediatricsGPT NEURIPS2024_fa5b423e, developed through systematic training for pediatric and general medical applications.
+These efforts show growing interest in adapting LLMs to medical reasoning and practice requirements.
+2.2 Evaluating Medical Knowledge in LLMs
+Numerous benchmarks evaluate LLM medical knowledge abacha2025medecbenchmarkmedicalerror.
+MedQA app11146421 is a multiple-choice QA dataset derived from the United States Medical Licensing Examinations that assesses clinical knowledge.
+HealthBench arora2025healthbenchevaluatinglargelanguage benchmarks realistic health conversations using physician-created rubrics.
+MedXpertQA zuo2025medxpertqabenchmarkingexpertlevelmedical evaluates expert-level medical knowledge through textual and multimodal tasks.
+Existing medical fact-checking benchmarks focus on narrow contexts.
+VeriFact Chung2025 verifies synthetic text against structured Electronic Health Records using retrieval-augmented generation (RAG) and LLM-as-a-Judge.
+MEDEC abacha2025medecbenchmarkmedicalerror provides error detection and correction benchmarks confined to clinical notes.
+By relying on synthetic data or single genres, these benchmarks fail to capture diverse real-world contexts, limiting their utility for assessing model generalization.
+2.3 Evaluating Factuality in LLMs
+Recent research has established benchmarks for factuality in LLMs.
+SimpleQA wei2024measuringshortformfactualitylarge measures short-form factuality using single-answer questions adversarially collected to challenge models such as GPT-4.
+OpenFactCheck iqbal-etal-2024-openfactcheck proposes a unified framework for benchmarking fact-checking across diverse sources, including free-form text and LLM-generated content.
+These benchmarks lack the domain-specific nuances required for high-stakes fields such as medicine.
+This highlights the need to bridge two gaps: the lack of domain-specific context in general factuality benchmarks and the lack of real-world diversity in medical fact-checking benchmarks.
+3 MedFact
+3.1 Data Collection
+The construction pipeline for MedFact is shown in Figure 1.
+We began with 27,116 texts secured through a copyright-compliant data-sharing agreement with a commercial partner.
+Sources include the partner’s internal encyclopedia primarily, supplemented with content from medical consultant platforms.
+As these texts are not publicly released, this mitigates the risk of data contamination.
+3.2 Data Curation
+Multi-criteria Filtering We defined a suite of seven leading LLMs: GPT-4o openai2024gpt4ocard, GPT-4.5 openai2024gpt45, o1 openai2024openaio1card, Gemini 2.5 Pro gemini, DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability, Claude 3.7 Sonnet claude37, and Doubao-Seed-1.6 doubao.
+This ensemble filtered the corpus via majority vote, discarding texts identified as overly simplistic (lacking medical knowledge), esoteric (highly specialized), or malformed (missing clear subject or context), achieving a Fleiss Kappa of 0.75 and pairwise consistency exceeding 80%.
+This process was refined through human-in-the-loop feedback inspired by wu2024autogen.
+In each round, experts reviewed 5% of filtered data.
+We used their annotations on misclassified items to construct more effective retrieval-augmented, few-shot prompts for subsequent rounds.
+Over three rounds, the acceptance rate decreased from 67.69% to 23.62%.
+The results achieved 96.40% agreement with expert judgments, reducing the number of candidates to 6,405 texts.
 Human Annotation The filtered corpus was then annotated by a team of three fully licensed medical professionals.
 The annotation for each text comprised: (a) a binary veracity judgment (correct/incorrect); (b) for incorrect texts, the precise error span; and (c) a suggested correction.
-Quality Enhancement and Refinement We then applied a series of automated techniques to enhance the dataset’s quality and difficulty, reducing the number of instances to 2,862. These steps included:
-  • Hard-case Mining: To increase the benchmark’s ability to differentiate among competent models, we evaluated all annotated instances with our model suite and removed those that all seven models classified correctly, thereby filtering for more challenging cases.
-  • Similarity Filtering: To promote topical and semantic diversity, we computed sentence embeddings for all texts using all-mpnet-base-v2 (Reimers and Gurevych 2019).
-For any pair of texts with a cosine similarity score exceeding a threshold of 0.9, we retained only one instance.
-  • Data Augmentation: To enhance linguistic diversity and ensure privacy, we prompted a randomly selected model from our suite to paraphrase and de-identify each text under two strict constraints: (a) preserving the original medical content; and (b) removing all Personally Identifiable Information (PII).
-Final Human Review In the final stage, the dataset underwent a comprehensive review by medical professionals.
-They verified that: (a) automated paraphrasing did not introduce errors or alter the original medical intent; (b) all PII was successfully removed; and (c) annotations remained accurate for the paraphrased texts.
-Any text failing these checks was either corrected or discarded.
-This process resulted in the final MedFact of 2,116 instances, which was balanced to contain an equal number of correct and incorrect texts.
-Figure 2: An example of fact-checking performed by Claude 3.7 Sonnet.
-The medical error is in red, and key steps in the model’s reasoning are in blue.
-The original Chinese text and reasoning are presented first, followed by their English translations.
-Figure 3: Distribution of the medical specialties in MedFact.
-Data Statistics & Analysis
+Quality Enhancement We applied automated techniques to enhance dataset quality and difficulty, reducing instances to 2,862:
+  • Hard-case Mining: We evaluated all instances with our model suite and removed those all models classified correctly, filtering for more challenging cases.
+  • Similarity Filtering: We computed sentence embeddings for all texts using all-mpnet-base-v2 reimers-2019-sentence-bert.
+For text pairs with cosine similarity exceeding 0.9, we retained only one instance.
+  • Data Augmentation: We prompted randomly selected models from our suite to paraphrase and de-identify each text under two constraints: (a) preserving original medical content; and (b) removing all Personally Identifiable Information (PII).
+Final Human Review Medical professionals conducted a review, verifying: (a) automated paraphrasing did not introduce errors or alter original medical intent; (b) all PII was removed; and (c) annotations remained accurate.
+Texts failing these checks were corrected or discarded.
+This process resulted in the final MedFact of 2,116 instances.
+3.3 Data Statistics & Analysis
 Overview MedFact contains 2,116 expert-annotated medical texts curated from real-world data.
@@
 It was constructed using a hybrid AI-human framework that combines large-scale multi-criteria filtering, hard-case mining, similarity filtering, and LLM-based augmentation with human expert review to ensure high data quality, diversity, and difficulty.
-Spanning 13 medical specialties, instances in MedFact are categorized by 8 fine-grained error types and multiple difficulty levels to facilitate a comprehensive evaluation of LLM fact-checking capabilities.
-Figure 2 presents a sample instance; additional examples are provided in the Appendix.
-Data Diversity MedFact covers 13 medical specialties, with a distribution led by Internal Medicine (1,156 instances), followed by Obstetrics and Gynecology (240), Surgery (231), and ten other fields.
-These are grouped into four high-level domains: Internal Medicine and Chronic Care; Surgical and Procedural Care; Specialized Clinical Fields; and Public and Mental Health (Figure 3).
-This distribution reflects a realistic range of medical topics, consistent with the distribution of 27,116 original medical texts, enabling robust evaluation of model performance across diverse clinical contexts.
-Beyond its topical breadth, MedFact exhibits significant stylistic diversity.
-The corpus is drawn from a variety of writing styles, including formal encyclopedia articles rich in professional terminology, popular science journalism, user-generated web content from both peer-to-peer and patient-physician Q&A, and fabricated or sensationalized misinformation.
-This diversity in topic and writing style ensures that MedFact mirrors the complexity of real-world medical information, facilitating a robust evaluation of an LLM’s generalizable fact-checking capabilities.
-Error Taxonomy Developed in consultation with medical professionals, our error taxonomy for MedFact distinguishes between two primary categories: medical and non-medical errors.
+Figure 2 presents an instance; additional examples are provided in the Appendix.
+Data Diversity MedFact covers 13 medical specialties, with a distribution led by Internal Medicine (1,156 instances), followed by Obstetrics and Gynecology (240), and eleven other fields (Figure 3).
+Beyond its topical breadth, MedFact exhibits stylistic diversity.
+The corpus is drawn from a variety of writing styles, including formal encyclopedia articles rich in professional terminology, popular science journalism, user-generated web content from peer-to-peer and patient-physician Q&A, and fabricated misinformation.
+This diversity ensures that MedFact mirrors the complexity of real-world medical information, facilitating a robust evaluation of an LLM’s generalizable fact-checking capabilities.
+Error Taxonomy Developed in consultation with professionals, our error taxonomy for MedFact distinguishes between two primary categories: medical and non-medical errors.
 Medical errors are subdivided into Conceptual, Terminological, Temporal, and Other Medical Errors, while non-medical errors are classified as Citation and Sourcing, Biased Content, General Factual, and Other Non-Medical Errors.
-Detailed definitions for each subcategory are available in the Appendix.
-As shown in Figure 1, the distribution across the 1,058 incorrect instances is heavily skewed toward medical errors (89.41%), with Conceptual errors being the most frequent subtype (52.65% of all incorrect instances).
-The prevalence of these medically grounded errors underscores that successful performance on this benchmark requires a deep understanding of medical principles.
-Difficulty Grading Instances in MedFact are stratified into five difficulty levels.
-This stratification is based on the aggregated success rate of our model suite on the error localization task (metrics are detailed in the Experimental Setup Section).
+As shown in Figure 1, the distribution across the 1,058 incorrect instances is skewed toward medical errors (89.41%).
+The prevalence of medically grounded errors underscores that successful performance on MedFact requires a deep understanding of medical principles.
+Difficulty Grading Instances in MedFact are stratified into five difficulty levels based on the aggregated success rate of our model suite on the error localization task (metrics are detailed in the Experiments Section).
 As shown in Figure 1, the resulting distribution across these levels is relatively balanced.
@@
 For each metric, the best score is in bold and the second-best is underlined.
-Human expert performance is provided for reference.
-Experiments
-Experimental Setup
-Models We evaluate a diverse set of 20 models on MedFact, comprising 8 open-source and 12 proprietary models:
-  • Open-source: II-Medical-8B (Internet 2025), HuatuoGPT-o1-7B (Chen et al.
-2024), DeepSeek-R1 (2025-05-28), DeepSeek-V3 (2025-03-24) (DeepSeek-AI 2025b), Qwen2.5-72B-Instruct (Team 2024), Qwen3-235B-A22B (Team 2025), DeepSeek-R1-Distill-Llama-70B (DeepSeek-AI 2025a), DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI 2025a).
-  • Proprietary: XiaoYi (2025-06-24) (Healthcare 2025), GPT-4o (2024-11-20), GPT-4.5, o1, o3 (OpenAI 2025b), Claude 3.7 Sonnet, Gemini 2.5 Flash (DeepMind 2025), Gemini 2.5 Pro, Grok-4 (xAI 2025), Qwen2.5-Max (Team 2024), ERNIE-X1 (Baidu 2025), Doubao-Seed-1.6 (ByteDance 2025).
-Among these models, XiaoYi, HuatuoGPT-o1-7B, and II-Medical-8B are specifically trained for the medical domain.
-We evaluate Doubao-Seed-1.6 in its standard version and thinking mode, listed as Doubao-Seed-1.6-thinking.
-We also establish a human expert baseline, calculated as the average performance of three medical professionals.
-Evaluation Strategies We evaluate all models using zero-shot and Chain-of-Thought (CoT) prompting (Wei et al.
-2022) in our main experiments.
-We also assess a subset of top-performing models using four advanced reasoning strategies, including recently proposed agent-based frameworks, which might be well-suited for fact-checking:
-  • MedPrompt (Nori et al. 2023): A composite strategy that combines self-generated CoT, retrieval-augmented few-shot examples, and self-consistency (Wang et al. 2023b) for single-model inference.
-  • Retrieval-Augmented Generation (RAG) (Lewis et al.
-2020): A technique that grounds model outputs by first retrieving relevant documents from an external knowledge corpus to use as context in the prompt.
-  • MAD (Liang et al. 2024): A multi-agent framework where agents engage in a multi-round debate to reach a consensus, thereby improving the final answer.
-  • MDAgents (Kim et al. 2024): A dynamic medical framework that adaptively engages multi-agent collaboration based on the perceived difficulty of the input.
-Evaluation Metrics We evaluate performance on two distinct tasks using Precision, Recall, and F1 score. The full implementation details are in the Appendix.
-  • Veracity Classification (VC): This binary classification task evaluates a model’s ability to identify a medical text as either correct or incorrect.
-The “incorrect” label serves as the positive class in calculating metrics.
-  • Error Localization (EL): For texts correctly classified as incorrect, this task evaluates the model’s ability to identify the span of text containing the error precisely.
-A prediction is considered a true positive only if the model localizes the correct error span.
-We employ GPT-4o as the automated judge for the EL task.
-Its reliability is confirmed by strong agreement with human experts (Cohen’s Kappa = 0.87) on a random 10% data sample, validating its use as a proxy for human evaluation.
+The human baseline is the average performance of three professionals.
+4 Experiments
+4.1 Experimental Setup
+Models We evaluate 20 models on MedFact:
+  • Open-source: II-Medical-8B 2025II-Medical-8B, HuatuoGPT-o1-7B chen2024huatuogpto1medicalcomplexreasoning, DeepSeek-R1 (2025-05-28), DeepSeek-V3 (2025-03-24) deepseekai2025deepseekv3technicalreport, Qwen2.5-72B-Instruct qwen2.5, Qwen3-235B-A22B qwen3technicalreport, DeepSeek-R1-Distill-Llama-70B deepseekai2025deepseekr1incentivizingreasoningcapability, DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability.
+  • Proprietary: XiaoYi (2025-06-24) xiaoyi, GPT-4o (2024-11-20), GPT-4.5, o1, o3 openai2025o3, Claude 3.7 Sonnet, Gemini 2.5 Flash gemini, Gemini 2.5 Pro, Grok-4 grok4, Qwen2.5-Max qwen2.5, ERNIE-X1 ernie2025technicalreport, Doubao-Seed-1.6 doubao.
+XiaoYi, HuatuoGPT-o1-7B, and II-Medical-8B are specifically trained for medical domains. We evaluate Doubao-Seed-1.6 in standard and thinking modes, listed as Doubao-Seed-1.6-thinking.
+Evaluation Strategies We evaluate models using zero-shot and Chain-of-Thought (CoT) prompting NEURIPS2022_9d560961, and assess top-performing models using four reasoning strategies:
+  • MedPrompt nori2023generalistfoundationmodelsoutcompete: A composite strategy that combines self-generated CoT, retrieval-augmented few-shot examples, and self-consistency wangself.
+  • RAG NEURIPS2020_6b493230: A technique that grounds model outputs by first retrieving relevant texts from an external knowledge corpus to use as context.
+  • MAD liang-etal-2024-encouraging: A multi-agent framework where agents engage in a multi-round debate to reach a consensus.
+  • MDAgents NEURIPS202490d1fc07: A medical framework adaptively engaging multi-agent collaboration based on the input’s difficulty.
+Evaluation Metrics We evaluate performance on two tasks using Precision, Recall, and F1 score:
+  • Veracity Classification (VC): Binary classification evaluating whether models identify medical texts as correct or incorrect. The “incorrect” label serves as the positive class.
+  • Error Localization (EL): For texts correctly classified as incorrect, this evaluates the model’s ability to identify the precise error span.
+A prediction is considered a true positive if the model correctly localizes the span.
+We use GPT-4o as a judge for EL. Its reliability is confirmed by strong agreement with human experts (Cohen’s Kappa = 0.87) on a random 10% sample.
 Model Veracity Classification Error Localization
@@
    +MDAgents 0.6038 0.8497 0.7059 0.5613 0.7136 0.6284
-Table 2: Performance of DeepSeek-R1 and XiaoYi across different strategies on MedFact. For each model, the best score is in bold and the second-best is underlined.
-Main Results
-Overall Performance Table 1 summarizes the performance of all evaluated models.
-We report the F1 score for VC and Precision, Recall, and F1 for EL.
-Among the models evaluated, proprietary models consistently outperform their open-source counterparts across both tasks and all prompting strategies.
-Doubao-Seed-1.6-thinking emerges as the top-performing general-purpose model, while XiaoYi leads among medically specialized models.
-A significant performance gap remains between the best-performing models and human experts.
-The top model F1 score on the EL task, 0.6858, achieved by XiaoYi with CoT, is still below the human baseline of 0.7012.
-This disparity underscores both the difficulty of MedFact and the substantial challenge for current LLMs.
-The full results are provided in the Appendix.
-A comparison between the tasks reveals a consistent performance gap, with all models performing worse on EL than on VC.
-For instance, the F1 scores for DeepSeek-V3 and Qwen2.5-Max decrease from 0.6947 and 0.7352 on VC to 0.5380 and 0.6116 on EL, respectively.
-This suggests that while models can often determine if a text contains an error, they struggle to pinpoint where it is.
-This difficulty often arises from a “correct-for-the-wrong-reason” phenomenon: a model may correctly classify a text as incorrect (a successful VC judgment) but do so by misidentifying a factually sound statement as the source of the error.
-Such behavior, rooted in a lack of deep medical knowledge, demonstrates that the EL task is a more stringent measure of a model’s fine-grained understanding and reveals fundamental deficiencies in its underlying medical knowledge.
-Impact of Advanced Reasoning Strategies The results of DeepSeek-R1 and XiaoYi with advanced reasoning strategies against a zero-shot baseline are presented in Table 10.
-The retrieval corpus for RAG consists of the 6,405 expert-annotated source texts from our data curation process, excluding any rewrites.
-Results show that RAG using either one or three retrieved references yields substantial performance gains on both the VC and EL tasks.
-This finding shows that providing models with direct access to external, domain-specific knowledge is a highly effective method for medical fact-checking.
-MedPrompt delivers modest improvements, demonstrating that performance can also be enhanced solely through prompting techniques such as few-shot learning and self-consistency.
-In contrast, the multi-agent frameworks (MAD and MDAgents) exhibit a consistent trade-off: they improve Recall at the expense of Precision, resulting in a stable F1 score for MAD and only a marginal improvement for MDAgents.
-We attribute this pattern to a phenomenon we term “over-criticism”.
-The collaborative process, which encourages the generation of multiple perspectives and error hypotheses, appears to increase the models’ propensity to flag errors where none exist, thereby elevating the false positive rate and lowering Precision.
-The slight performance edge of MDAgents may be attributed to its use of medical role-play, which could temper this critical tendency.
-This finding reveals a critical challenge: while these methods are designed to improve reasoning, they can inadvertently encourage an overly skeptical posture.
-This undermines reliability in medical domains, where correctly identifying accurate information is as crucial as flagging falsehoods.
+Table 2: Performance of DeepSeek-R1 and XiaoYi across different strategies on MedFact.
+4.2 Main Results
+Overall Performance Table 1 summarizes the evaluated model performance.
+Proprietary models consistently outperform open-source counterparts across both tasks and all prompting strategies.
+Doubao-Seed-1.6-thinking emerges as the top general-purpose model, while XiaoYi leads among medically specialized models.
+A significant performance gap remains between best-performing models and human experts.
+The top model F1 score on EL, 0.6858, achieved by XiaoYi with CoT, falls below the human baseline of 0.7012.
+Task comparison reveals that all models perform worse on EL than on VC.
+For instance, F1 scores for DeepSeek-V3 and Qwen2.5-Max decrease from 0.6947 and 0.7352 on VC to 0.5380 and 0.6116 on EL, respectively.
+This suggests that while models can often determine if text contains errors, they struggle to pinpoint the location.
+This difficulty often arises from a “correct-for-the-wrong-reason” phenomenon: models may correctly classify text as incorrect but misidentify factually sound statements as error sources.
+Such behavior, rooted in a lack of medical knowledge, demonstrates that EL is a more stringent measure of fine-grained understanding and reveals fundamental deficiencies in underlying medical knowledge.
+Advanced Reasoning Strategies Table 2 presents DeepSeek-R1 and XiaoYi results with advanced reasoning strategies against zero-shot baselines.
+The RAG corpus consists of 6,405 expert-annotated source texts from our data curation process.
+Results show RAG yields substantial performance gains on both VC and EL tasks.
+This demonstrates that providing models with direct access to external, domain-specific knowledge is highly effective for medical fact-checking.
+In contrast, multi-agent frameworks exhibit a trade-off: they improve Recall at the expense of Precision, resulting in stable F1 for MAD and marginal improvement for MDAgents.
+We attribute this to “over-criticism”: the collaborative process encourages generating multiple perspectives and error hypotheses, which increases models’ propensity to flag errors where none exist, thereby elevating false positive rates and lowering Precision.
+This reveals a challenge that, while these methods are designed to improve reasoning, they can encourage overly skeptical postures in medical fact-checking.
 Model F1 (VC) F1 (EL)
@@
 Table 3: Performance of different models with and without inference-time scaling techniques on MedFact.
-Analysis
-Inference-time Scaling
-Motivated by the superior performance of inference-time reasoning models such as DeepSeek-R1 and Doubao-Seed-1.6-thinking, over their models from the same series (DeepSeek-V3 and Doubao-Seed-1.6), we investigate the potential of inference-time scaling strategies.
-Specifically, we evaluate the impact of budget forcing (Muennighoff et al.
-2025).
-This technique modulates a model’s reasoning process at inference time by either prematurely terminating generation or prolonging it via “Wait” tokens to adjust the computational budget.
-We focus on two models adapted for this paradigm: s1.1-32B and m1-32B-1K.
-s1.1-32B is a general reasoning model fine-tuned from Qwen2.5-32B-Instruct on only 1,000 high-quality reasoning examples (Muennighoff et al.
-2025).
-m1-32B-1K is a specialized medical reasoning model fine-tuned from Qwen2.5-32B-Instruct, using 1,000 medical question-answering pairs (Huang et al.
-2025).
-Table 11 compares the performance of the baseline model (Qwen2.5-32B-Instruct) against s1.1-32B and m1-32B-1K under a CoT setting, evaluated both with and without budget forcing.
-The results indicate that s1.1-32B, fine-tuned on general-purpose data, suffers a substantial performance degradation on the EL task, which we attribute to its lack of domain-specific medical knowledge due to fine-tuning.
-In contrast, m1-32B-1K, with its medical fine-tuning, outperforms the base model on both tasks.
-However, applying budget forcing yields no performance gain and instead exacerbates the “over-criticism” phenomenon.
-The prolonged reasoning process causes models to hallucinate errors, a behavior analogous to the “over-thinking” observed in reasoning models (Ghosal et al.
-2025; Sui et al.
-2025).
-This suggests that simply extending a model’s reasoning time without providing new information can lead to counterproductive deliberation rather than improved accuracy.
-Ultimately, this finding reinforces the conclusion that for knowledge-intensive tasks like medical fact-checking, enhancing a model’s underlying medical knowledge is more critical than simply increasing its computational budget at inference.
+5 Analysis
+5.1 Inference-time Scaling
+We evaluate budget forcing (muennighoff2025s1simpletesttimescaling), which modulates model reasoning at inference time by either prematurely terminating generation or prolonging it via “Wait” tokens to adjust computational budget.
+We focus on two models adapted for this paradigm: s1.1-32B muennighoff2025s1simpletesttimescaling and m1-32B-1K.
+s1.1-32B is a general reasoning model fine-tuned from Qwen2.5-32B-Instruct on only 1,000 high-quality reasoning examples huang2025m1unleashpotentialtesttime.
+m1-32B-1K is a specialized medical reasoning model fine-tuned from Qwen2.5-32B-Instruct, using 1,000 medical question-answering pairs.
+Table 3 compares the performance of the baseline model against s1.1-32B and m1-32B-1K under a CoT setting.
+The results indicate that s1.1-32B, fine-tuned on general-purpose data, suffers a substantial performance degradation on the EL task, which we attribute to its lack of domain-specific medical knowledge.
+In contrast, m1-32B-1K, with its medical fine-tuning, outperforms the base model.
+Applying budget forcing yields no performance gain and instead exacerbates the “over-criticism” phenomenon.
+The prolonged reasoning process causes models to hallucinate errors, a behavior analogous to the “over-thinking” observed in reasoning models ghosal2025doesthinkinghelpunderstanding.
+This suggests that simply extending a model’s reasoning time without providing new information can lead to counterproductive deliberation, and reinforces the conclusion that for medical fact-checking tasks, enhancing a model’s underlying medical knowledge is more critical.
 Figure 4: Error distribution of DeepSeek-R1 and XiaoYi on the EL task (zero-shot).
-Error Analysis
-To better understand the performance gap between LLMs and human experts, and to diagnose why the EL task is particularly challenging, our human experts manually annotated and categorized the errors made by the top-performing models.
-As illustrated in Figure 4, the lack of medical knowledge constitutes the predominant category of model mistakes (76.09% of all errors for DeepSeek-R1), where models demonstrate insufficient understanding of medical principles, conflating distinct conditions or treatments, and relying on outdated information, which indicates that models often struggle with the nuanced, specialized knowledge required for accurate medical fact-checking.
-Fine-grained Performance
-We conduct a fine-grained analysis of model performance across different writing styles (Figure 9).
-For the VC task, models are more effective at identifying errors in fabricated medical misinformation than in texts from encyclopedias.
-This indicates that models can readily detect exaggerated falsehoods but struggle with the subtle, nuanced inaccuracies often present in professionally curated content.
-Notably, this performance gap diminishes for the more demanding EL task.
-This shift suggests that performance on localization is less influenced by stylistic cues and more fundamentally dependent on the model’s underlying medical knowledge.
-Consequently, the EL task serves as a more direct and robust measure of a model’s actual domain expertise.
 Model R-2 F1 R-L F1
@@
 Table 4: ROUGE scores of different models on the text completion task.
-Figure 5: Zero-shot performance of different models on MedFact across different writing styles.
-Data Contamination
+5.2 Error Analysis
+Human experts manually annotated and categorized top-performing model errors.
+As illustrated in Figure 4, lacking medical knowledge constitutes the predominant error category (76.09% of all errors for DeepSeek-R1), where models demonstrate insufficient understanding of medical principles, conflate distinct conditions or treatments, and rely on outdated information, indicating that models often struggle with nuanced, specialized knowledge required for accurate medical fact-checking.
+5.3 Fine-grained Performance
+We conduct fine-grained analysis of model performance across different writing styles (Figure 5).
+For VC, models are more effective at identifying errors in fabricated medical misinformation.
+This indicates that models can readily detect exaggerated falsehoods but struggle with subtle, nuanced inaccuracies often present in professionally curated content, and suggests that localization performance is less influenced by stylistic cues and more fundamentally dependent on underlying medical knowledge.
+This observation shows that EL serves as a more robust measure of medical expertise.
+5.4 Data Contamination
 To investigate potential data contamination within the pre-training corpora of the evaluated LLMs, we devised a text completion task.
@@
 As shown in Table 4, all evaluated models, even those that achieve state-of-the-art performance on our tasks, exhibit negligible ROUGE scores.
-This result strongly indicates that MedFact serves as a valid and previously unseen benchmark for these models.
-Conclusion
+This result indicates that MedFact serves as a valid and previously unseen benchmark for these models.
+Figure 5: Zero-shot performance of different models on MedFact across different writing styles.
+6 Conclusion
 We introduce MedFact, a new benchmark for evaluating the fact-checking abilities of LLMs in Chinese medical texts.
-Constructed via a meticulous hybrid AI-human pipeline, our dataset is diverse, realistic, and uncontaminated, allowing it to probe for nuanced medical understanding.
+Constructed via a hybrid AI-human pipeline, our dataset is diverse, realistic, and uncontaminated, allowing it to probe for nuanced medical understanding.
 Our comprehensive evaluation of 20 leading LLMs reveals a critical performance deficit compared to human experts, particularly on the fine-grained task of error localization.
@@
 The challenges highlighted underscore the substantial work required to develop LLMs that are not just knowledgeable but also factually reliable enough for safe deployment in medical applications.
-Appendix A Ethics Statement
-Human Annotation and Compensation All expert annotators who contributed to MedFact received fair compensation for their time and expertise at rates that meet or exceed the minimum wage in their respective regions.
-Privacy and Consent We employed a rigorous anonymization process to remove all Personally Identifiable Information (PII) and ensure privacy protection.
-Our use of this data is strictly for non-commercial research purposes and complies with the terms of service of source platforms under fair use principles.
-Risk of Misinformation The factually incorrect instances in MedFact could be exploited to train models that generate harmful content or to disseminate misinformation directly.
-We will distribute the dataset under a restrictive license that explicitly prohibits use in clinical or patient-facing applications.
-The dataset will be accompanied by comprehensive documentation that clearly warns users about the presence of factual errors and specifies that its intended purpose is exclusively for research and benchmarking applications.
-Bias and Representativeness MedFact focuses exclusively on Chinese medical texts.
-Consequently, the data, error distributions, and model findings may not generalize to other languages, cultures, or healthcare systems.
-We caution researchers against drawing overly broad conclusions from our findings and encourage the development of analogous benchmarks for other languages and cultural contexts.
+Limitations
+MedFact targets the Chinese language within China’s healthcare system, and performance may not generalize to other languages, cultures, or healthcare systems.
+This focus ensures high linguistic quality and cultural consistency, though future work could develop parallel benchmarks for cross-lingual comparison.
+For the EL task, we use GPT-4o as an automated judge, which enables scalable evaluation but has inherent limitations, including potential biases and inconsistencies.
+Future work could explore hybrid evaluation frameworks combining LLM judge scalability with expert validation for ambiguous or high-stakes cases.
+Additionally, MedFact captures medical knowledge at creation time, and as medical science evolves, some information may become outdated.
+We prioritize foundational medical knowledge to minimize this issue, but acknowledge that dynamic benchmark updates would be valuable in the long term.
+Ethical Considerations
+All expert annotators received fair compensation at or above minimum wage in their regions.
+We applied anonymization to remove all Personally Identifiable Information (PII), and data use is strictly for non-commercial research and complies with source platform terms under fair use principles.
+Factually incorrect instances could be exploited to train harmful content generators or disseminate misinformation, so we distribute the dataset under a restrictive license prohibiting clinical or patient-facing applications.
+The documentation clearly warns about factual errors and specifies that the dataset is intended exclusively for research and benchmarking purposes.
+Additionally, MedFact focuses exclusively on Chinese medical texts, and data, error distributions, and findings may not generalize to other languages, cultures, or healthcare systems.
+We caution against drawing overly broad conclusions and encourage the development of analogous benchmarks for other languages and contexts.
+Appendix A Detailed Related Works
+A.1 LLMs in Medical Domains
+The capabilities of Large Language Models (LLMs) across many disciplines have prompted the exploration of their applications in the medical domain huang2024comprehensivesurveyevaluatinglarge.
+Medprompt nori2023generalistfoundationmodelsoutcompete enhances LLMs’ performance on medical benchmarks by combining self-generated Chain-of-Thought (CoT) NEURIPS2022_9d560961, few-shot learning with dynamic example selection liu-etal-2022-makes, and self-consistency wangself.
+Other developments include HuaTuo wang2023huatuotuningllamamodel, a LLaMA-based model fine-tuned on Chinese medical instruction data; HuatuoGPT zhang-etal-2023-huatuogpt, a model specialized for medical consultation, fine-tuned on responses distilled from ChatGPT and real-world data; and PediatricsGPT NEURIPS2024_fa5b423e, a model developed via a systematic training pipeline for pediatric and general medical applications.
+These advances highlight a growing interest in adapting LLMs to medical reasoning and the specialized requirements of real-world practice.
+A.2 Benchmarking Medical Knowledge in LLMs
+Numerous benchmarks have been developed to evaluate the medical knowledge of LLMs nguyen-etal-2023-medredqa; NEURIPS2024_e15c4aff; abacha2025medecbenchmarkmedicalerror.
+MedQA app11146421 is a multiple-choice QA dataset derived from the United States Medical Licensing Examinations to assess the clinical knowledge of LLMs.
+Huatuo-26M wang-etal-2025-huatuo is a Chinese dataset comprising data from online consultation websites, encyclopedias, and information extracted from knowledge bases.
+More recently, HealthBench arora2025healthbenchevaluatinglargelanguage is a benchmark of realistic health conversations evaluated with custom, physician-created rubrics.
+MedXpertQA zuo2025medxpertqabenchmarkingexpertlevelmedical evaluates expert-level medical knowledge using both textual and multimodal tasks.
+While some benchmarks for medical fact-checking exist, they often focus on narrow contexts.
+VeriFact Chung2025 verifies synthetic text against structured Electronic Health Records using a retrieval-augmented generation (RAG) and LLM-as-a-Judge paradigm.
+MEDEC abacha2025medecbenchmarkmedicalerror provides a benchmark for error detection and correction, but is confined to clinical notes.
+Relying on synthetic data or a single text genre, these benchmarks fail to capture the diverse contexts of medical information encountered in the real world, thereby limiting their utility for assessing how well models generalize.
+Consequently, a gap remains for a benchmark that can robustly evaluate fact-checking across a wide array of realistic medical texts.
+A.3 Benchmarking Factuality in LLMs
+Recent research has produced several benchmarks to evaluate factuality in LLMs.
+For instance, HaluEval li-etal-2023-halueval focuses on hallucination detection through a large-scale collection of automatically generated and human-annotated examples.
+SimpleQA wei2024measuringshortformfactualitylarge measures short-form factuality using single-answer questions adversarially collected to challenge state-of-the-art models like GPT-4.
+More broadly, OpenFactCheck iqbal-etal-2024-openfactcheck proposes a unified framework for benchmarking fact-checking across diverse sources, including free-form text and LLM-generated content.
+While these benchmarks provide valuable tools for assessing general-domain factuality, they lack the domain-specific nuances required for high-stakes fields like medicine, where misinformation can have severe public health consequences 10.1001/jamapediatrics.2023.5282.
+Addressing this challenge requires a solution that bridges two critical gaps: the lack of domain-specific context in general factuality benchmarks and the lack of real-world diversity in existing medical fact-checking benchmarks.
+To this end, we introduce MedFact, a new benchmark designed to evaluate the fact-checking capabilities of LLMs on a comprehensive and diverse collection of realistic medical texts.
 Appendix B Dataset Construction & Statistics
-Multi-criteria Filtering
+B.1 Multi-criteria Filtering
 Across the three filtering rounds, the acceptance rates (ratio of passed instances to candidate pool size) were 67.69%, 37.00%, and 23.62%, respectively.
-The consistency between the model suite’s judgment and medical professionals increased progressively: 69.72%, 84.37%, and 96.40%, respectively, while the reference database size grew from 412 to 624 instances.
+The consistency between the model suite’s judgment and medical professionals increased progressively: 69.72%, 84.37%, and 96.40%, respectively, as the reference database size grew from 412 to 624 instances.
 Figure 10 presents one accepted example, and Figure 11 presents three rejected examples from the reference database during the multi-criteria filtering process, illustrating the filtering criteria: overly simplistic, esoteric, and malformed content.
-Error Taxonomy Definitions
+B.2 Error Taxonomy Definitions
 We define eight fine-grained error types in MedFact, organized into two primary categories: Medical Errors and Non-Medical Errors. The definitions for each subcategory are detailed below.
@@
 This includes failures to convey critical safety warnings, flawed clinical reasoning (e.g., broken causal chains between symptoms and diagnoses), or misrepresentation of risk factors.
-Non-Medical Errors Non-medical errors include factual inaccuracies that are not domain-specific to medicine.
+Non-Medical Errors Non-medical errors encompass factual inaccuracies that are not specific to the medical domain.
   • Citation and Sourcing Error: Fabrication or misrepresentation of sources, including citing non-existent literature, providing inaccurate citation details, or inventing data sources.
@@
   • Others (Non-medical): A residual category for other non-medical issues, including ethical or legal violations (e.g., privacy breaches, suggesting illegal acts), undisclosed commercial promotion, or severe linguistic incoherence that fundamentally obstructs comprehension.
-Writing Style Taxonomy Definitions
+B.3 Writing Style Taxonomy Definitions
 We categorize the writing styles of MedFact into four distinct types: encyclopedia-style articles, popular science journalism, user-generated web content, and fabricated misinformation:
@@
 Table 7: Distribution of samples in MedFact, categorized by accuracy score intervals.
-Data Statistics
+B.4 Data Statistics
 Tables 6 and 7 present the data distribution across different error types and difficulty levels, respectively. Difficulty levels are determined based on model performance.
-Data Examples
+B.5 Data Contamination
+To investigate potential data contamination within the pre-training corpora of the evaluated LLMs, we devised a text completion task.
+For each instance in MedFact, we prompted the models with the first half of the text and tasked them with generating the continuation.
+We then measured the textual overlap between the generated output and the ground-truth second half using ROUGE-F1 scores, where a high score would suggest that the model had memorized the content.
+As shown in Table 14, all evaluated models exhibit negligible ROUGE scores, which strongly indicates that MedFact serves as a previously unseen benchmark.
+B.6 Data Examples
 Figures 14 to 17 present examples for each medical error. Figures 18 to 21 present examples for each non-medical error. Figures 22 to 25 present examples for each writing style.
 Appendix C Implementation Details
-Prompt Templates
-Figures 12 to 13 present the prompts used for zero-shot and CoT prompting.
-All prompts were originally formulated in Chinese; we provide their corresponding English translations for clarity and reference.
-Model Versions
-We accessed proprietary models through API calls and deployed open-source models locally. Table 5 lists the model versions and full names used in our experiments.
-LLM-as-a-Judge
-For the evaluation of error localization tasks, we assessed the consistency agreement between different models and human experts.
+C.1 Prompt Templates
+Figures 12 to 13 present the prompts used for zero-shot and CoT prompting. All prompts were originally formulated in Chinese; we provide their corresponding English translations for reference.
+C.2 Model Versions
+We accessed proprietary models through API calls and deployed open-source models locally. Table 5 lists the model versions and full names.
+C.3 LLM-as-a-Judge
+For the evaluation of EL tasks, we assessed the agreement between different models and human experts.
 DeepSeek-R1, GPT-4o, o1, Claude 3.7 Sonnet, and Gemini 2.5 Pro achieved Cohen’s Kappa values of 0.800, 0.870, 0.870, 0.845, and 0.865, respectively, on a random 10% sample of the veracity classification task results from Doubao-Seed-1.6-thinking, XiaoYi, and Qwen-Max.
-Considering API costs and the evaluation performance, we selected GPT-4o for the LLM-as-a-judge evaluation.
-Appendix D Complete Experimental Results
-Full Result Tables
+Considering API costs and evaluation performance, we selected GPT-4o for the LLM-as-a-judge evaluation.
+Appendix D Experimental Results
+D.1 Full Result Tables
 We present the complete results of our main experiments in Tables 8 and 9 and Figures 6 and 7.
@@
 Table 11: Performance of different models with and without test-time scaling (budget forcing) on MedFact. The best results are bolded and the second-best are underlined.
+Model Precision Recall F1 Score
+DeepSeek-R1 0.5035 0.7580 0.6051
+   + RAG (top-3 from same sources) 0.6112 0.7713 0.6820
+   + RAG (top-3 from authoritative sources) 0.5557 0.7069 0.6223
+XiaoYi 0.6512 0.7023 0.6758
+   + RAG (top-3 from same sources) 0.6985 0.7117 0.7051
+   + RAG (top-3 from authoritative sources) 0.6861 0.6796 0.6828
+Table 12: Performance of different models with different sources for RAG on MedFact using zero-shot prompting. The best results are bolded and the second-best are underlined.
+Model Language Precision Recall F1 Score
+Gemini 2.5 Pro Chinese 0.4828 0.8752 0.6223
+English 0.5286 0.9317 0.6745
+Doubao-Seed-1.6-thinking Chinese 0.6307 0.7344 0.6786
+English 0.5952 0.8483 0.7000
+Table 13: Performance comparison between Chinese and English versions of MedFact using CoT prompting. The best results are bolded and the second-best are underlined.
 Model ROUGE-1 F1 ROUGE-2 F1 ROUGE-L F1 LCS Normalized Avg
@@
 Table 14: ROUGE scores of different models on the text completion task.
-Fine-grained Performance
-Figure 9 details the fine-grained model performance across various error types, medical specialties, and writing styles.
+D.2 Impact of RAG Reference Sources
+To evaluate how corpus selection affects medical fact-checking performance, we performed experiments comparing reference texts from authoritative sources (including peer-reviewed journals, clinical guidelines, and medical literature) with generic sources.
+Using the same retrieval pipeline, we assessed the impact of reference text relevance on fact-checking outcomes.
+As shown in Table 12, while precision improves slightly compared to zero-shot baselines when using authoritative sources, performance remains substantially lower than with highly relevant references.
+Crucially, recall decreases below the zero-shot baseline when using generic authoritative texts.
+We attribute this to relevance mismatch: irrelevant references severely degrade both precision and recall by biasing models toward accepting unverified content as correct, which shows a critical risk for industrial RAG deployments where reference quality varies.
+D.3 Cross-lingual Evaluation on English Translation
+To assess the impact of language on model performance, we evaluated an English-translated version of MedFact.
+We selected two models: Doubao-Seed-1.6-thinking and Gemini 2.5 Pro.
+Table 13 compares performance between Chinese and English versions using CoT prompting.
+Both models show F1 improvements on English: Doubao-Seed-1.6-thinking improves from 0.6786 to 0.7000, while Gemini gains more substantially from 0.6223 to 0.6745.
+However, both models consistently show decreased Precision and increased Recall in English.
+This indicates the over-criticism phenomenon persists and intensifies across languages.
+Models flag potential errors more aggressively in English, producing higher false positive rates despite improved F1 scores.
+These findings show language choice affects both overall performance and the precision-recall trade-off, with important implications for multilingual deployment in healthcare organizations.
+D.4 Fine-grained Performance
+We conduct fine-grained analysis of model performance across various error types, medical specialties, and writing styles (Figure 9).
+Across different writing styles, for VC, models are more effective at identifying errors in fabricated medical misinformation than in encyclopedia texts.
+This suggests that models can readily detect exaggerated falsehoods but struggle with subtle, nuanced inaccuracies that are often present in professionally curated content.
+This suggests that localization performance is less influenced by stylistic cues and depends more fundamentally on underlying medical knowledge.
+Consequently, EL serves as a more robust measure of actual domain expertise.
 Figure 9: Performance breakdown of different models on MedFact in the zero-shot setting.
 The radar charts illustrate model performance across three key dimensions: error type, medical specialty, and textual style.
-Data Contamination
+D.5 Data Contamination
 Table 14 presents the results of our data contamination analysis for the task completion experiment.
@@
 We further illustrate the “correct-for-the-wrong-reason” and “over-criticism” phenomena with case studies in Figures 42 to 46.
-Appendix F Limitations & Future Work
-Linguistic and Cultural Scope Our benchmark is developed exclusively for the Chinese language and is situated within the context of the Chinese healthcare system and common public discourse.
-The distribution and performance of LLMs may not generalize directly to other languages, cultures, or healthcare systems.
-This focus represents a deliberate methodological choice to ensure high linguistic quality and cultural consistency.
-Future work could involve creating parallel benchmarks in other languages to enable cross-lingual comparative studies.
-Single Error Assumption For clarity of annotation and evaluation, each incorrect text in MedFact contains only a single factuality error.
-This represents a simplifying assumption, as real-world misinformation often contains multiple, interwoven inaccuracies.
-Extending the benchmark to include instances with multiple, interacting errors would provide a more realistic and challenging evaluation of LLM fact-checking capabilities.
-Reliance on LLM-as-a-Judge For the EL task, we employ GPT-4o as an automated judge.
-While this enables scalable evaluation, it is subject to the inherent limitations of LLM judges, including potential biases and inconsistencies.
-Future research could explore hybrid evaluation frameworks that combine the scalability of LLM judges with targeted, expert-led validation for ambiguous or high-stakes cases.
-Static Knowledge MedFact represents a snapshot of medical knowledge at the time of its creation.
-As medical science evolves, some information may become outdated.
-Although we have prioritized foundational medical knowledge to minimize this issue, we acknowledge that establishing a process for dynamically updating the benchmark would constitute a valuable long-term objective.
 Figure 10: An example of accepted data with corresponding professional feedback.